### v1

# RoboSpatial: Teaching Spatial Understanding to 2D and 3D Vision-Language Models for Robotics

Abstract
Spatial understanding is a crucial capability for robots to make grounded decisions based on their environment. This foundational skill enables robots not only to perceive their surroundings but also to reason about and interact meaningfully within the world. In modern robotics, these capabilities are taken on by visual language models, and they face significant challenges when applied to spatial reasoning context due to their training data sources. These sources utilize general-purpose image datasets, and they often lack sophisticated spatial scene understanding capabilities. For example, the datasets do not address reference frame comprehension — spatial relationships require clear contextual understanding, whether from an ego-centric, object-centric, or world-centric perspective, which allow for effective real-world interaction. To address this issue, we introduce RoboSpatial, a large-scale spatial understanding dataset consisting of real indoor and tabletop scenes captured as 3D scans and egocentric images, annotated with rich spatial information relevant to robotics. The dataset includes 1M images, 5K 3D scans, and 3M annotated spatial relationships, with paired 2D egocentric images and 3D scans to make it both 2D and 3D ready. Our experiments show that models trained with RoboSpatial outperform baselines on downstream tasks such as spatial affordance prediction, spatial relationship prediction, and robotics manipulation.
11footnotetext: Corresponding author: song.1855@osu.edu. This work was conducted during the first author’s internship at NVIDIA.
1 Introduction
Figure 1: An illustration of a model trained on RoboSpatial being used to solve a manipulation task using spatial reasoning. We show that the trained model can enable human-aligned spatial reasoning in the correct reference frame. Furthermore, our chosen spatial relationships enable models to do task grounding, planning, and successful detection.
The introduction of vision-and-language models (VLMs) in modern robotics has helped reconcile general computer vision with robotics control [59]. We have seen these models trained directly on robotics manipulation data [10] which allows robots to directly consume image and task description inputs and act in the real-world. Other works have used generic VLMs to help describe robotics scenes to accomplish specific tasks such as cleaning a table top [16]. We have also seen grasping augmented with text-based description [49]. Moreover, large language models (LLMs) are also important to modern robotics, they have been used extensively to write robot code from user prompts [44, 29] or on high-level planning [1, 45, 21].
Dataset 3D scans Embodied Frames Compatibility Domain #Scans #Images #Spatial QAs
EmbSpatial-Bench [13] Indoor 277    2K    4K
Visual Spatial [33] MSCOCO 0 10K 10K
SpatialRGPT-Bench [9] Indoor, AV 0 1.4K 1.4K
BLINK-Spatial [17] Generic 0   286   286
What’s up [26] Generic 0    5K   10K
Spatial-MM [43] Generic 0 2.3K 2.3K
RoboSpatial Indoor, tabletop 5K    1M    3M
Table 1: Comparison with other spatial reasoning datasets that include object-centric spatial relationships.
VLMs can recognize objects, classify scenes, and even provide general descriptions that capture high-level attributes. However, despite significant recent advancements, VLMs [40, 34, 30] still fall short in spatial understanding [56, 26, 43]. They struggle with tasks that require interpreting nuanced spatial relationships between objects, such as describing where one object is in relation to another or determining the best location to place an item within a specific condition. For example, while a model might accurately describe a “bowl on the table,” it lacks the ability to reason about where on the table the bowl is, where it should go to ensure accessibility or stability, or how it might fit among other objects. Furthermore, a critical limitation of existing VLM training datasets is their inability to capture reference frame understanding — the way we interpret spatial relationships changes drastically depending on whether we’re viewing from a first-person perspective, focusing on specific objects, or observing the entire scene, all of which are essential for real-world interactions. These limitations underscore the gap between surface-level scene description and the deeper spatial comprehension necessary for intuitive interaction.
To address this gap, various works have proposed methods to teach VLMs about spatial reasoning. Although some recent works have explored training VLMs for spatial understanding, they stop short of meeting the requirements for robotics applications. For example, SpatialVLM [8] and SpatialRGPT [9] train VLMs to answer questions about distances and spatial relationships between objects, advancing spatial understanding at a conceptual level. However, these models are trained on datasets comprised of images from the internet, with annotations generated by perception models. As a result, they struggle to generalize to images captured by robot cameras that often lack identifiable cues for an absolute scale. RoboPoint [57] takes a different approach by training a VLM to produce grounded 2D coordinates, pinpointing object locations or free space within a scene. Yet, due to its reliance on synthetic data, the model lacks an understanding of real-world constraints, such as the amount of space needed to place specific objects or the suitability of surfaces for various items. It is also not able to infer object-centric reference frames, such as identifying what is in front of a car. Most recently, Molmo [12], a suite of vision-language models designed for object centric image-space pointing, shows promise for bridging AI and physical environments. However, our findings reveal limitations in its understanding of real-world constraints, e.g., failing to predict whether the gray bowl can fit in front of the car in Figure 1.
Our hypothesis is that one of the main bottlenecks limiting the effectiveness of VLMs in robotics is the limited training data itself, which is clearly highlighted by Table 1. Therefore, we introduce RoboSpatial, a dataset designed to teach spatial understanding to VLMs with robotic applications in mind. Our approach leverages annotated indoor scene and tabletop RGBD datasets, transforming them into targeted question-answer pairs that probe specific spatial reasoning skills relevant for robotics. We categorize the questions into three types: (1) Spatial configuration, focusing on object-object spatial relationships; (2) Spatial context, addressing object-space spatial relationships; and (3) Spatial compatibility, exploring object-object compatibility. Each question type serves a distinct purpose: (1) The spatial configuration questions focus on binary assessments, asking whether certain spatial relationships exist between pairs of objects, such as “Is the mug to the left of the laptop?” (2) Spatial context questions are designed as point-prediction tasks, challenging the model to determine appropriate locations within free space where an object can be placed, such as “Where on the table can I put the plate?” Finally, (3) the spatial compatibility questions also use a binary format to address compatibility between objects, such as “Can the chair be placed in front of the table?”
To enhance the model’s ability to interpret spatial instructions from different perspectives, each question-answer pair in RoboSpatial is posed from three distinct reference perspectives/frames: (a) Ego-centric from the observer’s perspective at the camera pose, (b) Object-centric based on a reference frame attached to the focal object, and (c) World-centric grounded in a global world frame. This multi-frame approach enables models to handle complex spatial instructions more flexibly, preparing them to better generalize to dynamic robotic contexts. Applying our methodology to existing indoor scene and tabletop datasets, we generate both a comprehensive training dataset and a benchmark for spatial question answering in robotics. RoboSpatial contains around 1M images, 5k 3D scans, and 3M annotated spatial relationships, with paired 2D egocentric images and 3D scans to make it both 2D and 3D ready, see Figure 1 for an example taken from our dataset.
To validate the effectiveness of our dataset, we conduct comprehensive experiments with multiple state-of-the-art (SOTA) 2D and 3D VLMs. Results show that training VLMs on RoboSpatial significantly enhances their spatial reasoning capabilities, with models trained on our dataset outperforming existing models on our test split. To further demonstrate the robustness of RoboSpatial, we assess the trained VLMs on three additional downstream robotic tasks: RoboSpatial-Home, , a high-quality manually annotated indoor scene question-answering dataset, BLINK-Spatial [17], and real-world robot manipulation. These tasks directly evaluate the models’ ability to apply spatial understanding in practical scenarios, such as rearranging objects or responding to context-specific spatial questions within an indoor environment. Additionally, we test our models on an external benchmark, BLINK-Spatial [17], to measure generalization to unseen spatial reasoning tasks beyond our dataset. Leveraging RoboSpatial’s 3D-ready design, we also compare the performance of 2D and 3D VLMs directly. These results emphasize the strengths of RoboSpatial in advancing spatial understanding and reinforce the benefits of 3D-based training for enhancing the spatial reasoning capabilities of VLMs in robotic applications.
Our contributions are threefold:
  • A new training dataset and benchmark, RoboSpatial and RoboSpatial-Home, comprising images and 3D scans paired with spatial questions and answers, designed with unique features including multiple reference frames, object-object spatial relationships, object-space relationships, and object compatibility. We plan to publicly release our data and code for generating the dataset from 3D annotated scenes.
  • Training of VLMs on RoboSpatial, resulting in enhanced spatial reasoning capabilities that significantly outperform prior VLMs on natural language-specified robot manipulation tasks and indoor scene question answering.
  • Comprehensive experiments assessing spatial reasoning capabilities in both 2D and 3D VLMs, comparing the difference between SOTA VLMs in real-world spatial tasks.
2 Related Work
VLMs for Robotics. Vision-language models (VLMs) have emerged as pivotal tools in robotics, enabling systems to interpret and act upon complex visual and textual information. By integrating visual perception with language understanding, VLMs facilitate more intuitive human-robot interactions and enhance autonomous decision-making capabilities. Recent advancements have demonstrated the potential of VLMs in various robotic applications. For instance, vision-and-language action models (VLAs) [27, 4, 39] enable robots to interpret and execute complex instructions and output executable robot actions. Additionally, VLMs like GPT-4v [40] have been utilized for high-level task planning [51], allowing robots to generate detailed action sequences from natural language instructions. Furthermore, VLMs have been used for keypoint prediction [22, 16, 38], error analysis [14], grasp pose prediction [19]. Despite these advancements, integrating VLMs [5, 9, 57] into robotic systems presents challenges. One significant hurdle is the need for precise spatial reasoning to navigate and manipulate objects effectively. While VLMs excel in understanding and generating language, their ability to comprehend and reason about spatial relationships in dynamic environments remains limited [56, 55, 53]. Therefore, RoboSpatial aims to tackle this gap by presenting a large scale pretraining and evaluation setup for teaching spatial understanding to VLM for robotics.
Figure 2: Overview of the RoboSpatial dataset. We automatically generate spatial relationship annotations from existing datasets with 3D point clouds, egocentric images, and 3D bounding box annotations. We create question/answer pairs covering three classes of spatial relationships, three spatial reference frames, and both binary (yes/no) and numeric (e.g. 2D image points) answers. From 1M images and 1.4K scans, we generate over 2M spatial question/answer pairs.
Spatial Understanding with VLMs. Spatial understanding has been implicitly and explicitly part of various vision and question answering tasks [17, 3, 24, 46, 2, 28, 25, 23]. While many benchmarks and methods have been proposed, they often come with limitations: some focus exclusively on simulations [48] or generic images [33, 41, 9, 8, 17, 26, 43], others are difficult to evaluate [48, 13], rely on complete 3D scans [58, 37, 36, 32], do not consider reference frames [58, 37, 36, 32, 8, 9, 17], or use data generated by LLMs that may lack sufficient accuracy [32]. Furthermore, they often fail to address actionable, robotics-relevant spatial relationships such as spatial compatibility and context [13, 17, 54, 43, 26].
Inspired by prior works on spatial reasoning [33, 26]—where the impact of reference frames and spatial configurations was explored in generic images [31, 23]—we extend spatial understanding to a robotics-specific context with actionable spatial relationships such as spatial compatibility and spatial context. Our aim is to enable direct application to robotic workflows, such as task planning and verification.
To achieve this, we have developed and are planning to open-source a large-scale 2D/3D ready pretraining dataset, an automated data annotation pipeline, and trained models. We further show how our dataset can be used to teach spatial reasoning to a suite of vision-language models (VLMs) in in-domain and out-of-domain spatial reasoning datasets. We hope these resources lower the barrier to entry for exploring spatial understanding tailored to robotics.
3 Approach
In this section, we begin by explaining our selection of three spatial relationships: spatial configuration, spatial context, and spatial compatibility. Next, we provide an overview of the data generation pipeline used to construct RoboSpatial. An overview of the RoboSpatial is provided in Figure 2.
3.1 Spatial Relationships
Our dataset is organized around three core spatial relationships that we believe to address the essential aspects of spatial reasoning for robotic tasks: object configuration, object context, and object compatibility. Configuration enables robots to understand and interpret the relative positioning of objects, which is crucial for directing navigation, manipulation, and interaction within complex environments. Context allows robots to assess the relationship between objects and their surrounding space, facilitating the identification of empty or occupied areas and aiding in efficient path planning and obstacle avoidance. Compatibility focuses on whether objects can coexist or interact without conflict in a given space, which is vital for object placement, assembly, and operational safety. Together, these spatial relationships provide a more nuanced and practical framework for robotic applications than metrics like distance—which is hard to normalize across different scales, environments, and tasks—thereby enabling robots to perform complex tasks with greater reliability.
3.2 Dataset Generation
The goal of the data construction pipeline is to generate a high-accuracy spatial relationship dataset with minimal human intervention by use of carefully constructed heuristics. Our data generation process takes as input a source scene dataset $\mathcal{D}_{s}$ that contains RGB views, camera poses, and object annotations in form of text labels and oriented 3D bounding boxes. The output is a dataset $\mathcal{D}$ where each datum $d_{i}=\langle I_{i},q_{i},a_{i},l_{i}\rangle$ consists of an image $I_{i}$, a question $q_{i}$, an answer $a_{i}$, and a reference frame label $l_{i}\in\{\textit{ego},\textit{world},\textit{object}\}$. Each question belongs to one of the three core spatial reasoning categories or a fourth category for object grounding (Table 2).
Our data generation process consists of two stages, (1) Spatial Relation Extraction, and (2) Question-Answer Generation.
3.2.1 Spatial Relation Extraction
In the spatial relation extraction stage, we analyze the dataset to discover a set of spatial relations between objects or points in free space. Each spatial relation can be described as $\langle I_{i},a_{i},t_{i},r_{i},l_{i}\rangle$, where $I_{i}$ is an image, $a_{i}$ is an anchor object, $t_{i}$ is a target object or a target free-space point, $s_{i}$ is size measure based on an object to be manipulated, $r_{i}\in\{\textit{left},\textit{right},\textit{above},\textit{below},\textit{front},\textit{behind}\}$ is a spatial relation preposition, and $l_{i}\in\{\textit{ego},\textit{world},\textit{object}\}$ is a reference frame label.
We use oriented 3D bounding box annotations to automatically generate spatial relationships. Since the 3D bounding boxes contain information about orientation (which direction the object is facing) and location in the world (including elevation), we then utilize a heuristic to automatically extract spatial relationships between objects and space.
Spatial Configuration: Spatial configuration relationships involve determining whether one object is in a certain spatial direction relative to an anchor object. We first choose an anchor object and generate spatial relationships with all the objects that appear in the image. For each of the three reference frames, we select the direction indicated by the spatial relation. For example, if the spatial relation is (in front, object frame), we only select objects that are in the forward direction from the oriented bounding box. To avoid ambiguity, we only consider objects that appear once in the current frame and generate relationships between all object pairs in the image. Answers for this relationship are binary (True/False).
Spatial Context: Spatial context relationship refers to a task where the model is required to determine points (x,y coordinates in the camera frame, also known as image space) relative to the spatial relationship with the anchor object. One example of such a relationship is identifying points that are in front of another object. Specifically, we construct a top-down map of the environment using the annotated 3D bounding boxes and randomly sample points in empty areas that are a set distance from the object. This distance could be based on the size of another object that is to be manipulated and placed in relation to the object. We apply the same reference frame logic as described for the spatial configuration heuristic. Answers to this relationship are numeric, i.e., a list of 2D coordinates that fit the criteria.
Spatial Compatibility: Spatial compatibility is an extension of the spatial context relationship where the model needs to reason about whether a referenced object can fit in a region indicated relative to the anchor object. We use the same top-down map constructed for spatial context generation and simulate placing an object in the map using its bounding box size. We apply the same reference frame logic to define the area we should simulate placing an object. Answers for this relationship are binary (True/False).
Question Types Category Answer type
Spatial Relations Spatial Configuration True/False
Spatial Context 2D points
Spatial Compatibility True/False
Grounding Object Referring 2D bounding box
Table 2: Question Taxonomy.
3.2.2 Question-Answer Generation
Based on the generated relationships, we construct three types of question-answer pairs corresponding to each spatial relationship type: spatial configuration, spatial context, and spatial compatibility. To ensure that the model reasons with vision rather than relying on language commonsense, we structure the questions and answers to be template-based and unambiguous. We use a template: {object/space} {relationship} {anchor object} {reference frame}, where the spatial relationship and reference frame is defined in  Sec. 3.2.1. We maintain a balanced dataset with respect to spatial relationships to avoid introducing bias into the model. Furthermore, we use the 2D bounding box to create an auxiliary object-referring dataset to teach models what the object is, in order to avoid errors arising from a lack of object grounding. Using our pipeline we generated 3M spatial relationships, which is an order of magnitude larger than other spatial reasoning datasets shown in Table 1. Figure 2 shows an overview of the generated question types.
4 Experiments
4.1 Implementation
We apply our data generation process to three scene datasets and two tabletop datasets. The scene datasets include ScanNet [11], Matterport3D [7], and 3RScan [52]. The object tabletop datasets are HOPE [50] and GraspNet-1B [15]. We retrieve the 3D bounding box annotations and embodied images from EmbodiedScan [54]. This mix of datasets provides coverage for diverse indoor scenes featuring large objects for navigation, and smaller objects for learning priors for robot manipulation. In total, we generate a large scale dataset for spatial reasoning that contains around 3M spatial QA pairs with 5K 3D scans and 1M images. Statistics of these datasets are described in Table 3.
Dataset Type Splits Images QA pairs
Indoor Train 4914 scans 880K images 2.6M
Validation 50 scans 1K images 3K
Test 50 scans 1K images 3K
Tabletop Train 150 scenes 75K images 217K
Validation 40 scenes 3K images 3K
Test 77 scenes 3K images 3K
Table 3: Dataset splits for indoor and tabletop dataset. Detailed data statistics are in the Appendix.
Figure 3: In-domain and out-of-domain results for RoboSpatial-trained models. Two models are shown: SL (SpaceLLaVA [8]) and RP (RoboPoint [57]), where the S- prefix and FT indicates RoboSpatial-trained models. For Yes/No questions, the green checkmark indicates the correct answer. For spatial context questions, GT indicates the correct answer. For out-of-domain, all images except for the top right are from RoboSpatial-Home.
4.1.1 Trained 2D/3D VLMs
2D VLMs.  For our dataset, we want to compare a suite of VLMs, including ones fine-tuned for spatial understanding and ones that are not. We chose two base VLMs: VILA-1.5-8B [30] and LLaVA-NeXT-8B [34]; three specialized VLMs: SpaceLLaVA-13B, a community implementation of SpatialVLM [8]; RoboPoint-13B [57], a model fine-tuned on synthetic datasets to predict points in empty areas referenced by an object; and Molmo-7B [12], a new family of VLMs specializing in pointing and counting We also include a closed-source model, GPT-4o [40]. We omit models such as SpatialRGPT [9] due to its reliance on external object mask generation as input, which omits the important aspect of object grounding.
3D VLMs. Extending VLMs to the 3D domain introduces unique challenges due to the inherent complexity of 3D data. The scarcity of large-scale, high-quality 3D datasets hinders the development of end-to-end 3D VLMs. As such, there are few open source VLMs for 3D understanding that reason from 3D inputs. Among the available models, we selected: 3D-LLM [18], which lifts multi-view images into a 3D point cloud, and LEO [20], which directly takes as input 3D object point clouds by relying on a pre-processing object segmentation step. These models allow us to evaluate spatial understanding in 3D by processing data either through reconstructed 3D representations or directly from point clouds. By incorporating both approaches, we aim to assess the effectiveness of 3D VLMs in handling spatial information from different types of 3D data inputs.
Fine-tuning. Our evaluation includes both “zero-shot” evaluation of each VLM, plus evaluation after fine-tuning on RoboSpatial for open-source VLMs. See the Appendix for further details.
4.1.2 Spatial Understanding Evaluation
We evaluate our method in answering 3000 questions (1000 questions per spatial relationship) from our test set, and measure how often the model produces the correct answer. The evaluation includes binary yes/no questions, and questions requiring numeric responses. For yes/no questions, we report the success rate in producing the correct response. For numeric questions, we report the rate at which the produced numeric answer is within the convex hull of a set of points generated by our spatial relation extraction method. Note that this is potentially an overly strict metric, and therefore the reported performance on numeric questions reflects a lower-bound of the true performance. Table 4 shows the main results.
4.1.3 Cross-Dataset Generalization Evaluation
Next, we evaluate the generalization capability of our method by testing it across different scene types—specifically, both indoor and tabletop scenes—to control for any bias in the annotations of the underlying datasets that make up our benchmark. We train on data derived from subsets of the datasets corresponding to one scene type (either indoor or tabletop) and test on held-out datasets from the other scene type, representing unseen environments. We expect that even when training on a subset of datasets, the performance on unseen scene types will improve if our method generalizes well. The results of this cross-dataset evaluation are shown in Table 5.
4.1.4 Out-of-Domain Evaluation
To test the out-of-domain transferability of RoboSpatial-trained models, we evaluate them on RoboSpatial-Home, our manually annotated indoor scene dataset, and the validation spatial reasoning split of the Blink [17] benchmark.
RoboSpatial-Home is a collection of 50 crowd-sourced indoor RGBD images captured using a handheld iPhone camera equipped with a built-in depth sensor. Each image is manually annotated with three spatial relationships, resulting in a total of 150 questions across 50 images. We designed RoboSpatial-Home to test the RoboSpatial-trained model’s capability to generalize to unseen objects in real life cluttered scenes.
Blink [17] is a visual reasoning benchmark designed to evaluate the reasoning capabilities of vision-language models (VLMs) using visual input alone. Blink was chosen to assess the generalization ability of RoboSpatial-trained models on unseen spatial relationships such as “next to,” “touching,” and “on top.” Unlike benchmarks such as What’s Up [26], which include spatial relationships similar to those in RoboSpatial, Blink [17] provides a more rigorous test of out-of-domain generalization. The validation spatial reasoning split of Blink consists of 143 binary yes/no questions about spatial configurations.
The results of the out-of-domain generalization evaluations are presented in Table 6 and visualized in Fig. 3.
Model Indoor Tabletop Average
Configuration Context Compatibility Configuration Context Compatibility Indoor Tabletop Total
Open-source VLMs
2D VLMs
VILA [30] 54.7 18.3 56.3 45.1 13.2 53.8 43.1 37.4 40.2
+RoboSpatial 71.4 $\uparrow$ 45.9 $\uparrow$ 77.2 $\uparrow$ 71.8 $\uparrow$ 43.7 $\uparrow$ 73.3 $\uparrow$ 64.8 $\uparrow$ 62.9 $\uparrow$ 63.9 $\uparrow$
LLaVA-Next [34] 48.9 12.5 32.7 48.3 8.4 30.9 31.4 29.2 30.3
+RoboSpatial 69.3 $\uparrow$ 41.3 $\uparrow$ 70.5 $\uparrow$ 70.7 $\uparrow$ 44.8 $\uparrow$ 66.1 $\uparrow$ 60.4 $\uparrow$ 60.5 $\uparrow$ 60.5 $\uparrow$
SpaceLLaVA [8] 52.6 15.3 49.0 66.5 12.2 60.1 38.9 46.2 43.6
+RoboSpatial 76.0 $\uparrow$ 50.7 $\uparrow$ 76.6 $\uparrow$ 74.9 $\uparrow$ 46.4 $\uparrow$ 70.5 $\uparrow$ 67.8 $\uparrow$ 63.6 $\uparrow$ 65.7 $\uparrow$
RoboPoint [57] 39.0 41.4 38.3 37.9 31.6 45.2 39.6 38.2 38.9
+RoboSpatial 72.2 $\uparrow$ 68.9 $\uparrow$ 72.1 $\uparrow$ 70.3 $\uparrow$ 61.7 $\uparrow$ 78.4 $\uparrow$ 71.0 $\uparrow$ 70.1 $\uparrow$ 70.6 $\uparrow$
3D VLMs
3D-LLM [18] 54.5 8.1 53.6 59.2 10.6 57.4 37.6 42.4 40.0
+RoboSpatial 76.3 $\uparrow$ 35.4 $\uparrow$ 77.5 $\uparrow$ 76.2 $\uparrow$ 46.8 $\uparrow$ 75.0 $\uparrow$ 63.1 $\uparrow$ 66.0 $\uparrow$ 64.6 $\uparrow$
LEO [20] 56.1 11.3 58.3 60.8 11.1 59.3 41.9 43.7 42.8
+RoboSpatial 80.2 $\uparrow$ 56.7 $\uparrow$ 82.5 $\uparrow$ 78.1 $\uparrow$ 55.2 $\uparrow$ 78.9 $\uparrow$ 73.1 $\uparrow$ 70.7 $\uparrow$ 71.9 $\uparrow$
Not available for fine-tuning
2D VLMs
Molmo [12] 40.6 48.2 60.0 61.5 35.8 54.6 49.6 50.6 50.1
GPT-4o [40] 63.5 25.1 59.4 62.3 27.9 66.8 49.3 52.3 50.8
Table 4: Results of existing 2D/3D VLMs on a held-out test split of images and scans. All methods, for all tasks, perform better ($\uparrow$) when fine-tuned on our RoboSpatial dataset. The best result for each column is bolded.
Indoor $\rightarrow$ Tabletop Tabletop $\rightarrow$ Indoor
RoboPoint [57] 38.7 38.2
+RoboSpatial 48.9 $\uparrow$ 51.3 $\uparrow$
LEO [20] 41.9 43.7
+RoboSpatial 47.2 $\uparrow$ 54.5 $\uparrow$
[…]
Model RoboSpatial-Home BLINK-Spatial
Localization Affordance Compatibility Accuracy
Open-source VLMs
2D VLMs
VILA [30] 53.3 12.0 52.0 72.7
+RoboSpatial 62.0 $\uparrow$ 32.0 $\uparrow$ 58.0 $\uparrow$ 79.7 $\uparrow$
LLaVA-Next [34] 48.0 9.3 37.3 71.3
+RoboSpatial 58.0 $\uparrow$ 24.0 $\uparrow$ 44.0 $\uparrow$ 79.0 $\uparrow$
SpaceLLaVA [8] 60.0 16.0 49.3 76.2
+RoboSpatial 68.7 $\uparrow$ 38.0 $\uparrow$ 56.0 $\uparrow$ 81.8 $\uparrow$
RoboPoint [57] 43.3 41.3 36.0 63.6
+RoboSpatial 50.0 $\uparrow$ 54.0 $\uparrow$ 48.0 $\uparrow$ 70.6 $\uparrow$
3D VLMs
3D-LLM [18] 40.0 8.0 46.0 N/A
+RoboSpatial 48.0 $\uparrow$ 36.0 $\uparrow$ 52.7 $\uparrow$ N/A
LEO [20] 50.7 10.0 48.0 N/A
+RoboSpatial 64.0 $\uparrow$ 40.0 $\uparrow$ 60.0 $\uparrow$ N/A
Not available for fine-tuning
Molmo [12] 44.7 38.0 58.0 67.1
GPT-4o [40] 64.0 25.3 56.7 76.2
Table 6: Results on an out-of-domain test split comparing prior art VLMs. The results show improved ($\uparrow$) spatial understanding capabilities on similar domains. Bolded number is the best result for the column.
4.2 Results
Our experiments on the held-out test split in Table 4, out-of-domain splits in Table 6, and real-world robot experiments in Table 7 clearly demonstrate the power of RoboSpatial in teaching spatial reasoning to VLMs where training on RoboSpatial improved spatial understanding across all spatial reasoning benchmarks. We further answer the following detailed questions about  RoboSpatial-trained models:
How well does RoboSpatial training generalize to unseen spatial relationships? While RoboSpatial contains templated question-answer pairs and a fixed set of spatial relationship prepositions, Table 6 shows that RoboSpatial training does generalize to other spatial relationship prepositions present in Blink [17], such as “under,” “next to,” and “far away.” We highlight that RoboSpatial encompasses all six principal directions along the x, y, and z axes in 3D space, and that generalizing to novel relationships involves mapping each of these directions to one or more spatial relationship prepositions, a task at which LLMs excel. For example, “on top of” and “on” refer to the “above” direction in a world-centric frame, while “under” corresponds to “below.” Furthermore, we observe that prepositions such as “next to” and “beside” imply that objects need to be in close proximity to one another. Since RoboSpatial contains spatial context questions that teach the model to generate points close to a given object, this implicitly teaches what it means to be “close” to another object, enabling RoboSpatial-trained models to understand concepts like “next to.” However, RoboSpatial training does not seem to generalize well to questions involving human-centric perspectives, since we focus solely on objects in our dataset for robotics use cases.
Do RoboSpatial-trained models understand nuanced perspectives? When humans refer to spatial relationships, we often imply a specific frame of reference. For example, “in front of the car” typically refers to the car’s front hood. Therefore, questions in RoboSpatial-Home omit explicitly specifying the reference frame to test the model’s ability to align its understanding with the questioner’s intended reference frame. Results in Table 6 show that models trained with RoboSpatial can automatically infer the reference frame, with Figure 3 showing examples like “Is the frame in front of the nightstand.”
Are 3D VLMs better at learning spatial relationships than 2D VLMs? While our results seem to imply that 3D VLMs outperform their 2D counterparts in learning spatial relationships, we point out that 3D-LLM [18] and LEO [20] are trained with indoor scan datasets, making the comparison between 2D and 3D models not entirely fair. Nevertheless, we hypothesize that 3D models have an advantage in spatial understanding due to their ability to incorporate depth information. Therefore, we developed RoboSpatial to be both 2D and 3D ready to support further research in this area.
Model Success Rate (%)
Open-source
LLaVA-Next [34] 23.7
RoboSpatial 52.6 $\uparrow$
RoboPoint [57] 44.7
RoboSpatial 46.2 $\uparrow$
Not available for fine-tuning
Molmo [12] 43.8
GPT-4o [40] 46.9
Table 7: Success rate comparison between robotics systems powered by zero-shot VLMs and VLMs fine-tuned on RoboSpatial.
4.3 Real Robot Experiments
Figure 4: Robotics experiments: the red dot shows the model output (if not present, the model failed to provide a valid point in the image); green dots are used to show when a model outputs multiple points. The robot motion generator, cuRobo [47], is used to grasp the item referenced by the generated point. The spatial- prefix indicates model trained with RoboSpatial.
We designed a comprehensive suite of tasks that require spatial reasoning and object manipulation. We use a diverse set of objects, including various shapes (e.g., cubes, cylinders, boxes), colors (red, orange, blue, green), and everyday items (food items, toys) as shown in Figure 4. Our goal in selecting these objects is not to confound the methods with difficult-to-identify items but to make identification as simple as possible with no ambiguity. For each scene configuration, we asked two yes/no questions, followed by asking a pick-and-place question. We note that picking and placing were treated as independent tasks for evaluations. We conducted over 200 model queries during our experiments, where details of questions and scenes used in the experiments are in the Appendix. For robot experiments, we used a Kinova Jaco robot [6] paired with a ZED2 camera for RGB(D) perception and employ cuRobo [47] for collision-free motion planner.
We evaluate the following VLMs: LLaVA-NeXT [34] and RoboPoint [57], both with and without RoboSpatial training; and Molmo [12] and GPT-4o [40] as baselines. Table 7 and Fig 4 presents the results of the real-world robot experiments. Our results show that for generic VLM such as LLaVa-Next [34], training on RoboSpatial significantly improved its spatial understanding capabilities in robot manipulation tasks. Furthermore, RoboSpatial-trained models were more likely to align predictions with the implied reference frame, as demonstrated in the “place in front of the pony” task, where only RoboSpatial-trained models’ predictions aligned correctly in front of the pony’s head. Similarly, in the “place in front of the orange juice box” task, RoboSpatial-trained models could identify an appropriate distance to place the object relative to its size. In contrast, for RoboPoint, the predicted point was too far to the object, making it difficult to place the object correctly. However, we note that GPT-4o performs comparably with RoboSpatial-trained RoboPoint. We attribute this to GPT-4o being able to understand much more diverse language inputs, unlike RoboSpatial, which was trained on templated language.
Here, we show that our dataset enhances the performance of 2D VLMs in robotics tasks. In our robot setup, 2D VLMs project predicted points from pixel space into 3D space; however, even a 2-pixel error can translate into a displacement of 5 to 10 cm in 3D space, leading to task failure. Interesting future direction is exploring how perspective impacts 2D point selection in the context of VLMs. Another interesting future direction is to enable 3D VLMs to capably support partial 3D point cloud observations, unlike the models we currently train, which require complete colored 3D point clouds, so that they can be effectively utilized in real-world robotics scenarios.
5 Conclusion
We introduce RoboSpatial and RoboSpatial-Home, a large-scale 2D/3D spatial understanding training and evaluation dataset tailored for robotics. Experimental results show that models trained with RoboSpatial are able to understand spatial relationships, generalize to unseen relationships, and infer nuanced reference frames, making them applicable in a wide range of tasks that require spatial understanding. We further demonstrate the real-world applicability of RoboSpatial with robot experiments. Furthermore, our automatic data generation pipeline can be used to extend the dataset whenever needed. We show that RoboSpatial has the potential to serve as a foundation for broader applications in robotic scenarios that require spatial understanding.
Acknowledgement
The authors would like to thank Youngsun Wi, Hyunho Ahn, Minjae Bae, and Hojin Yoo for providing images in the RoboSpatial-Home dataset. Our research is supported in part by ARL W911NF2220144 and resources from the Ohio Supercomputer Center. These research findings were partially derived using the Matterport dataset, the use of which is governed by its terms of service: https://kaldir.vc.in.tum.de/matterport/MP_TOS.pdf.
Appendices
In this supplementary material, we present additional details and clarifications that are omitted in the main text due to space constraints.
Appendix A Dataset Details
A.1 Dataset Statistics
We provide the full dataset statistics in Tab. 8. For all training, we use only 900,000 spatial relationships, equally divided across all datasets, due to computational constraints. We further experiment on the effect of data scaling on Tab. 9 and explain the results. Notably, HOPE [50] and GraspNet-1B [15] contain similar tabletop images captured from different perspectives, resulting in lower dataset diversity for the tabletop environment. We plan to enhance the diversity of our dataset by incorporating additional tabletop datasets.
A.2 Dataset Generation Details
Frame Generation. We explain how answers are generated for each frame. Each question type has three answers from ego-centric, object-centric, and world-centric perspectives. Each frame may share the same answer since not all frames lead to unique answers. For our answer generation, we used the following logic:
  • Ego-centric: The default perspective from the ego-centric view.
  • Object-centric: Using the oriented bounding box directions, we determine the front of the object. With this information, we assign front, behind, left, and right. Above and below remain the same as in the ego-centric perspective.
  • World-centric: Using the z-coordinate of the oriented bounding box, we modify the above and below relationships to reflect whether an object is above or below another object with respect to elevation.
Compatibility Generation. For compatibility, we construct a top-down map as shown in Fig. 5. Using the top-down map and the top-down 2D bounding box of the object to be placed, we determine if there exists an empty space within a threshold distance (i.e., 1% of the object’s longest width or length). If the object meets this condition, it is deemed compatible to be fitted.
Category Dataset Split Scans Images Configuration Q Context Q Compatibility Q
Indoor Matterport3D [7] Train 1859 scans 236243 298439 298439 298439
Validation 5 scans 100 100 100 100
Test 5 scans 100 100 100 100
ScanNet [11] Train 1512 scans 278402 298439 298439 298439
Validation 12 scans 1200 500 500 500
Test 12 scans 1200 500 500 500
3RScan [52] Train 1543 scans 365355 298439 298439 298439
Validation 9 scans 900 400 400 400
Test 9 scans 900 400 400 400
Tabletop HOPE [50] Train 50 scenes 50000 36317 36317 36317
Validation 10 scenes 50 500 500 500
Test 47 scenes 235 500 500 500
GraspNet-1B [15] Train 100 scenes 25500 36317 36317 36317
Validation 30 scenes 120 500 500 500
Test 30 scenes 120 500 500 500
Table 8: Full dataset statistics for indoor and tabletop datasets.
Figure 5: An example of generated top-down map of the image from 3D bounding boxes.
Appendix B Implementation Details
B.1 Model Training
We further explain the training details for all 2D and 3D VLMs trained on RoboSpatial. For all models, we perform instruction tuning using the model weights from public repositories. All training is done using 8 Nvidia H100 GPUs, with the training time between 20 and 40 hours.
VILA [30] We initialize our model from Efficient-Large-Model/Llama-3-VILA1.5-8B on Hugging Face. We use the fine-tuning script from the VILA GitHub repository to train our model using the default hyperparameters.
LLaVA-NeXT [34] We initialize our model from lmms-lab/llama3-llava-next-8b on Hugging Face. We use the LLaVA-Next fine-tuning script from the LLaVA-Next repository using the default hyperparameters.
SpaceLLaVA [8] As official code and weights for SpatialVLM [8] is not released, we use a community implementation which is endorsed by SpatialVLM [8] authors. We initialize our model from remyxai/SpaceLLaVA from Hugging Face. We use LLaVA-1.5 finetuning script from LLaVa [35] repository using the default hyperparameters.
RoboPoint [57] We initialize our model from wentao-yuan/robopoint-v1-vicuna-v1.5-13b on Hugging Face. We use the fine-tuning script provided in the RoboPoint [57] GitHub repository to train our model using the default hyperparameters.
3D-LLM [18] We initialize our model using the pretrain_blip2_sam_flant5xl_v2.pth checkpoint downloaded from the official GitHub repository. Since the model requires preprocessing of multiview images, we follow the author’s pipeline to process multiview images from our environments. Because the model does not accept image input, we append the following text in front of the question to ensure the model understands the perspective from which the question is being asked: “I am facing anchor object.” We use the default hyperparameters and train the model for 20 epochs per the author’s guidelines. We choose the best model based on validation accuracy.
LEO [20] We initialize our model from the sft_noact.pth checkpoint downloaded from the official GitHub repository.
Since LEO supports dual image and 3D point cloud input, we input both of them and modify the question as in 3D-LLM. We use the default hyperparameters and train the model for 10 epochs per the author’s guidelines, and choose the best model based on validation accuracy.
We could not fine-tune Molmo [12] from allenai/Molmo-7B-D-0924 or GPT-4o [40] from the gpt-4o-2024-08-06 API due to the unavailability of the fine-tuning script at the time of this work, thus we use them as a zero-shot baselines.
Annotation Size 100K 300K 900k (Default) 1.8M 3M (Full)
LLaVa-Next [34] 38.1 46.7 60.5 65.8 72.4
Table 9: Results of scaling experiment on LLaVa-Next [34] with varied spatial relationship annotations. Average accuracy on held-out test set is reported.
B.2 Robot Setup
For picking, we find which object the point maps to using SAM 2 [42] and execute our picking behavior on that object. For placing, we simply compute the 3D coordinate based on the depth value at that pixel and place the object at that coordinate. There were no failures due to cuRobo [47] failing. The experiments were purposely designed to consist of behaviors that our robot system can handle in order to avoid introducing irrelevant factors. The picking behavior consists of computing a top-down grasp pose and reaching it with cuRobo [47]. To compute the grasp pose:
  • 1.
    We estimate the major axis of the object’s point cloud in top-down view using PCA.
  • 2.
    The grasp orientation is orthogonal to the major axis.
  • 3.
    The grasp height is based on the highest point in the object’s point cloud minus an offset of 3cm. This heuristic ensures the system can grip long objects.
The placing behavior is the same as picking, except that an area within 5cm of the placement coordinate is used as the point cloud for estimating orientation and height, and a vertical height offset is added to account for the height at which the object was picked.
Model Indoor Tabletop Average
Ego-centric Object-centric World-centric Ego-centric Object-centric World-centric Indoor Tabletop Total
Open-source VLMs
2D VLMs
VILA [30] 55.9 40.5 32.9 43.6 39.7 28.9 43.1 37.4 40.2
+RoboSpatial 74.3$\uparrow$ 57.8 $\uparrow$ 62.3 $\uparrow$ 70.3 $\uparrow$ 58.1 $\uparrow$ 60.3 $\uparrow$ 64.8 $\uparrow$ 62.9 $\uparrow$ 63.9 $\uparrow$
LLaVA-Next [34] 35.2 24.3 34.7 36.4 28.5 22.7 31.4 29.2 30.3
+RoboSpatial 75.4 $\uparrow$ 54.1 $\uparrow$ 68.8 $\uparrow$ 67.9 $\uparrow$ 54.7 $\uparrow$ 58.9 $\uparrow$ 60.4 $\uparrow$ 60.5 $\uparrow$ 60.5 $\uparrow$
SpaceLLaVA [8] 40.6 36.0 30.1 52.3 32.8 53.5 38.9 46.2 43.6
+RoboSpatial 78.5 $\uparrow$ 60.6 $\uparrow$ 64.3 $\uparrow$ 73.0 $\uparrow$ 49.5 $\uparrow$ 68.3 $\uparrow$ 67.8 $\uparrow$ 63.6 $\uparrow$ 65.7 $\uparrow$
RoboPoint [57] 41.9 36.2 40.7 46.2 30.5 37.9 39.6 38.2 38.9
+RoboSpatial 76.4 $\uparrow$ 58.3 $\uparrow$ 78.3 $\uparrow$ 76.7 $\uparrow$ 62.6 $\uparrow$ 71.0 $\uparrow$ 71.0 $\uparrow$ 70.1 $\uparrow$ 70.6 $\uparrow$
3D VLMs
3D-LLM [18] 28.9 38.3 45.6 38.9 35.7 52.6 37.6 42.4 40.0
+RoboSpatial 60.7 $\uparrow$ 52.1 $\uparrow$ 76.5 $\uparrow$ 57.9 $\uparrow$ 62.8 $\uparrow$ 77.3 $\uparrow$ 63.1 $\uparrow$ 66.0 $\uparrow$ 64.6 $\uparrow$
LEO [20] 46.9 30.6 48.2 41.4 34.3 55.4 41.9 43.7 42.8
+RoboSpatial 68.1 $\uparrow$ 71.6 $\uparrow$ 79.6 $\uparrow$ 71.4 $\uparrow$ 60.2 $\uparrow$ 80.5 $\uparrow$ 73.1 $\uparrow$ 70.7 $\uparrow$ 71.9 $\uparrow$
Not available for fine-tuning
2D VLMs
Molmo [12] 50.4 50.8 47.6 64.4 33.6 53.8 49.6 50.6 50.1
GPT-4o [40] 52.9 38.7 56.3 62.5 30.7 63.7 49.3 52.3 50.8
Table 10: Results of per frame accuracy of existing 2D/3D VLMs on a held-out test split of images and scans. All methods, for all tasks, perform better ($\uparrow$) when fine-tuned on our RoboSpatial dataset. The best result for each column is bolded.
Appendix C Full Results
C.1 Data Scaling
In Tab. 9, we experiment with scaling the number of annotations while keeping images fixed. We found that even though the number of images stays consistent, increasing the number of annotations can improve performance. For future work, we plan to apply our data generation pipeline to a diverse set of indoor and tabletop environments to further improve the performance of our models.
C.2 Accuracy Per Frame
We show the results per frame in Tab. 10 for our out-of-domain test set. From the results, we can see a distinct difference between 2D and 3D VLMs in understanding the world-centric frame before training with RoboSpatial. Baseline 2D VLMs have trouble understanding the world-centric frame, which involves understanding elevation, while 3D VLMs comparatively excel at it. Furthermore, we can see that since baseline 3D VLMs are trained on point clouds without information of perspective, their accuracy in ego-centric and object-centric frames is lower. However, with RoboSpatial training, we were able to teach the 3D VLMs to think in a certain frame, thus considerably improving their performance on ego-centric and object-centric frames. However, we hypothesize that, due to their design—specifically, the lack of a means to visually inject perspective information since they require complete 3D point clouds—3D VLMs still lag behind 2D VLMs on ego-centric and object-centric frames.
C.3 Robot Experiments
We present additional results from our robot experiments in Fig. 6. We observe that models trained with RoboSpatial consistently outperform baseline models in most cases, even though the prompt is not optimized for RoboSpatial-trained models. This demonstrates that the power of VLMs enables templated language to generalize to language unseen during training while maintaining spatial understanding capabilities. However, even with RoboSpatial training, the models struggle with understanding stacked items, indicating a need for further data augmentation with diverse layouts. In a few cases, RoboSpatial training adversely affects performance, especially with RoboPoint [57]. We hypothesize that mixing the dataset with RoboPoint training data and RoboSpatial training data may lead to unforeseen side effects, particularly in grounding objects. Nevertheless, we demonstrate that RoboSpatial training enhances VLM’s spatial understanding in real-life robotics experiments, even with freeform language.
Question: pick lone object
GPT-4o [40] ×
Molmo [12]
LLaVa-Next [34] ×
S-LLaVa-Next [34]
RoboPoint [57] ×
S-RoboPoint [57]
Question: Is there room to slot the pancake mix in the middle of the row of boxes
GPT-4o [40]
Molmo [12]
LLaVa-Next [34]
S-LLaVa-Next [34]
RoboPoint [57] ×
S-RoboPoint [57]
Question: Is there space in the white container for the orange juice box
LLaVa-Next [34] ×
S-LLaVa-Next [34]
RoboPoint [57] ×
S-RoboPoint [57] ×
Molmo [12] ×
GPT-4o [40]
Question: pick object behind the middle container
LLaVa-Next [34] ×
S-LLaVa-Next [34]
RoboPoint [57]
S-RoboPoint [57] ×
Molmo [12] ×
GPT-4o [40] ×
Question: place object in container behind popcorn
LLaVa-Next [34] ×
S-LLaVa-Next [34]
RoboPoint [57]
S-RoboPoint [57]
Molmo [12] ×
GPT-4o [40] ×
Question: alphabet soup fit in the purple box
LLaVa-Next [34]
S-LLaVa-Next [34] ×
RoboPoint [57]
S-RoboPoint [57]
Molmo [12] ×
GPT-4o [40]
Question: pick shortest object
LLaVa-Next [34] ×
S-LLaVa-Next [34]
RoboPoint [57]
S-RoboPoint [57]
Molmo [12]
GPT-4o [40]
Question: place the object inside the smallest box
LLaVa-Next [34] ×
S-LLaVa-Next [34]
RoboPoint [57]
S-RoboPoint [57]
Molmo [12]
GPT-4o [40] ×
Question: can the robot directly pick the red orange peaches can without disturbing other objects?
LLaVa-Next [34]
S-LLaVa-Next [34]
RoboPoint [57] ×
S-RoboPoint [57] ×
Molmo [12]
GPT-4o [40]
Question: can the macaroni and cheese be placed on top of cheez-it without touching other objects?
LLaVa-Next [34] ×
S-LLaVa-Next [34] ×
RoboPoint [57]
S-RoboPoint [57]
Molmo [12] ×
GPT-4o [40]
Question: place on the object to the left of macaroni and cheese
GPT-4o [40] ×
Molmo [12]
LLaVa-Next [34] ×
S-LLaVa-Next [34]
RoboPoint [57]
S-RoboPoint [57]
Question: is there an object that is not in a stack?
GPT-4o [40]
Molmo [12]
LLaVa-Next [34]
S-LLaVa-Next [34]
RoboPoint [57]
S-RoboPoint [57]
Question: is there space to place one of the cans on the cheez-it box?
GPT-4o [40] ×
Molmo [12] ×
LLaVa-Next [34] ×
S-LLaVa-Next [34] ×
RoboPoint [57] ×
S-RoboPoint [57] ×
Question: pick the highest object on the stack of two objects
GPT-4o [40] ×
Molmo [12] ×
LLaVa-Next [34] ×
S-LLaVa-Next [34] ×
RoboPoint [57] ×
S-RoboPoint [57] ×
Figure 6: Additional robot experiments. A green check mark indicates that the model answered correctly. The S- prefix denotes a model trained with RoboSpatial. The questions are purposely not cleaned to reflect realistic language inputs.
Appendix D More Qualitative Examples
Fig. 7 and Fig. 8 present additional qualitative comparisons between models trained on RoboSpatial. Our findings demonstrate that models trained on RoboSpatial consistently exhibit spatial understanding in the challenging RoboSpatial-Home dataset, even outperforming closed models like GPT-4o [40]. However, we observed that object grounding is a crucial prerequisite for spatial understanding; the improvement is often hindered by the model’s inability to ground objects in cluttered scenes, where GPT-4o performs more effectively. Additionally, in Fig. 8, we show that the RoboSpatial-trained model successfully generalizes to unseen spatial relationships in Blink-Spatial [17], including those involving distance, such as ”touching.”
Figure 7: Qualitative results on RoboSpatial-Home. For spatial context questions, we omit the baseline models which fail to generate a point inside the image. Furthermore, the first predict point for each model is shown in the image.
Figure 8: Qualitative results on Blink-Spatial [17]. RoboSpatial-trained model can generalize to unseen spatial relationships.

### v2

@@
 Abstract
-Spatial understanding is a crucial capability for robots to make grounded decisions based on their environment.
-This foundational skill enables robots not only to perceive their surroundings but also to reason about and interact meaningfully within the world.
-In modern robotics, these capabilities are taken on by visual language models, and they face significant challenges when applied to spatial reasoning context due to their training data sources.
-These sources utilize general-purpose image datasets, and they often lack sophisticated spatial scene understanding capabilities.
-For example, the datasets do not address reference frame comprehension — spatial relationships require clear contextual understanding, whether from an ego-centric, object-centric, or world-centric perspective, which allow for effective real-world interaction.
-To address this issue, we introduce RoboSpatial, a large-scale spatial understanding dataset consisting of real indoor and tabletop scenes captured as 3D scans and egocentric images, annotated with rich spatial information relevant to robotics.
-The dataset includes 1M images, 5K 3D scans, and 3M annotated spatial relationships, with paired 2D egocentric images and 3D scans to make it both 2D and 3D ready.
+Spatial understanding is a crucial capability that enables robots to perceive their surroundings, reason about their environment, and interact with it meaningfully.
+In modern robotics, these capabilities are increasingly provided by vision-language models.
+However, these models face significant challenges in spatial reasoning tasks, as their training data are based on general-purpose image datasets that often lack sophisticated spatial understanding.
+For example, datasets frequently do not capture reference frame comprehension, yet effective spatial reasoning requires understanding whether to reason from ego-, world-, or object-centric perspectives.
+To address this issue, we introduce RoboSpatial, a large-scale dataset for spatial understanding in robotics.
+It consists of real indoor and tabletop scenes, captured as 3D scans and egocentric images, and annotated with rich spatial information relevant to robotics.
+The dataset includes 1M images, 5k 3D scans, and 3M annotated spatial relationships, and the pairing of 2D egocentric images with 3D scans makes it both 2D- and 3D- ready.
 Our experiments show that models trained with RoboSpatial outperform baselines on downstream tasks such as spatial affordance prediction, spatial relationship prediction, and robotics manipulation.
-11footnotetext: Corresponding author: song.1855@osu.edu. This work was conducted during the first author’s internship at NVIDIA.
+11footnotetext: Corresponding author: song.1855@osu.edu. This work was partly done during the first author’s internship at NVIDIA.
 1 Introduction
-Figure 1: An illustration of a model trained on RoboSpatial being used to solve a manipulation task using spatial reasoning.
-We show that the trained model can enable human-aligned spatial reasoning in the correct reference frame.
-Furthermore, our chosen spatial relationships enable models to do task grounding, planning, and successful detection.
-The introduction of vision-and-language models (VLMs) in modern robotics has helped reconcile general computer vision with robotics control [59].
-We have seen these models trained directly on robotics manipulation data [10] which allows robots to directly consume image and task description inputs and act in the real-world.
-Other works have used generic VLMs to help describe robotics scenes to accomplish specific tasks such as cleaning a table top [16].
-We have also seen grasping augmented with text-based description [49].
-Moreover, large language models (LLMs) are also important to modern robotics, they have been used extensively to write robot code from user prompts [44, 29] or on high-level planning [1, 45, 21].
-Dataset 3D scans Embodied Frames Compatibility Domain #Scans #Images #Spatial QAs
+Figure 1: RoboSpatial dataset facilitates 3D spatial reasoning for robotic manipulation.
+This illustration demonstrates how a model trained on RoboSpatial enables human-aligned spatial reasoning within the correct reference frame, supporting task grounding, planning, and detection for manipulation tasks.
+The rise of vision-language models (VLMs) has opened new opportunities for agents to interpret and act upon the visual world using natural language.
+VLMs have been adopted across a range of embodied settings, notably in robotics and augmented reality (AR).
+In robotics, they have enabled grounded scene understanding [13, 66], manipulation [7], and policy code generation [48, 29], while in AR, they support tasks like object labeling [50], action recognition [17, 16], and temporal grounding [22].
+Dataset 3D scans Embodied Ref. frames Compatibility Domain #Scans #Images #Spatial QAs
 EmbSpatial-Bench [10] Indoor 277    2k    4k
@@
 VLMs can recognize objects, classify scenes, and even provide general descriptions that capture high-level attributes.
-However, despite significant recent advancements, VLMs [40, 34, 30] still fall short in spatial understanding [56, 26, 43].
+However, despite significant recent advancements, VLMs [42, 35, 30] still fall short in spatial understanding [61, 26, 47, 38].
 They struggle with tasks that require interpreting nuanced spatial relationships between objects, such as describing where one object is in relation to another or determining the best location to place an item within a specific condition.
 For example, while a model might accurately describe a “bowl on the table,” it lacks the ability to reason about where on the table the bowl is, where it should go to ensure accessibility or stability, or how it might fit among other objects.
-Furthermore, a critical limitation of existing VLM training datasets is their inability to capture reference frame understanding — the way we interpret spatial relationships changes drastically depending on whether we’re viewing from a first-person perspective, focusing on specific objects, or observing the entire scene, all of which are essential for real-world interactions.
-These limitations underscore the gap between surface-level scene description and the deeper spatial comprehension necessary for intuitive interaction.
-To address this gap, various works have proposed methods to teach VLMs about spatial reasoning.
-Although some recent works have explored training VLMs for spatial understanding, they stop short of meeting the requirements for robotics applications.
+Furthermore, a critical limitation of existing VLM training datasets is their inability to capture reference frame understanding (ref.
+frame) — the way we interpret spatial relationships changes drastically depending on whether we’re viewing from a first-person perspective, focusing on specific objects, or observing the entire scene, all of which are essential for real-world interactions.
+These limitations highlight an ongoing challenge: bridging the gap between surface-level scene description and the deeper spatial comprehension necessary for intuitive interaction.
+Several recent efforts aim to address this by explicitly training VLMs on spatial reasoning tasks, yet many fall short of the demands posed by embodied or robotics settings.
 For example, SpatialVLM [5] and SpatialRGPT [6] train VLMs to answer questions about distances and spatial relationships between objects, advancing spatial understanding at a conceptual level.
 However, these models are trained on datasets comprised of images from the internet, with annotations generated by perception models.
-As a result, they struggle to generalize to images captured by robot cameras that often lack identifiable cues for an absolute scale.
-RoboPoint [57] takes a different approach by training a VLM to produce grounded 2D coordinates, pinpointing object locations or free space within a scene.
-Yet, due to its reliance on synthetic data, the model lacks an understanding of real-world constraints, such as the amount of space needed to place specific objects or the suitability of surfaces for various items.
-It is also not able to infer object-centric reference frames, such as identifying what is in front of a car.
-Most recently, Molmo [12], a suite of vision-language models designed for object centric image-space pointing, shows promise for bridging AI and physical environments.
-However, our findings reveal limitations in its understanding of real-world constraints, e.g., failing to predict whether the gray bowl can fit in front of the car in Figure 1.
-Our hypothesis is that one of the main bottlenecks limiting the effectiveness of VLMs in robotics is the limited training data itself, which is clearly highlighted by Table 1.
-Therefore, we introduce RoboSpatial, a dataset designed to teach spatial understanding to VLMs with robotic applications in mind.
-Our approach leverages annotated indoor scene and tabletop RGBD datasets, transforming them into targeted question-answer pairs that probe specific spatial reasoning skills relevant for robotics.
-We categorize the questions into three types: (1) Spatial configuration, focusing on object-object spatial relationships; (2) Spatial context, addressing object-space spatial relationships; and (3) Spatial compatibility, exploring object-object compatibility.
-Each question type serves a distinct purpose: (1) The spatial configuration questions focus on binary assessments, asking whether certain spatial relationships exist between pairs of objects, such as “Is the mug to the left of the laptop?” (2) Spatial context questions are designed as point-prediction tasks, challenging the model to determine appropriate locations within free space where an object can be placed, such as “Where on the table can I put the plate?” Finally, (3) the spatial compatibility questions also use a binary format to address compatibility between objects, such as “Can the chair be placed in front of the table?”
-To enhance the model’s ability to interpret spatial instructions from different perspectives, each question-answer pair in RoboSpatial is posed from three distinct reference perspectives/frames: (a) Ego-centric from the observer’s perspective at the camera pose, (b) Object-centric based on a reference frame attached to the focal object, and (c) World-centric grounded in a global world frame.
+As a result, they struggle to generalize to embodied images—those captured by robot cameras within real-world environments, which often lack identifiable cues for an absolute scale.
+Pointing models, such as RoboPoint [62] and more recently Molmo [9], take a different approach by training VLMs to produce grounded 2D coordinates that pinpoint object locations or free space within a scene.
+However, these models lack understanding of real-world constraints, such as inferring object-centric reference frames for perspective-invariant reasoning, or accounting for the space required to place various objects.
+This results, for example, in failing to predict whether the gray bowl can fit in front of the car in Figure 1.
+This paper hypothesizes that a primary bottleneck limiting the effectiveness of VLMs in robotics is the scarcity of suitable training data, as highlighted by Table 1.
+To address this, we introduce RoboSpatial, a dataset designed specifically to facilitate spatial understanding in VLMs for robotic applications.
+The proposed approach leverages annotated indoor scene and tabletop RGBD datasets, transforming them into targeted question-answer pairs designed to probe spatial reasoning skills critical for robotics.
+We categorize the questions into three types, each serving a distinct purpose.
+Spatial context focuses on identifying empty space or support surfaces in the environment that can accommodate other objects.
+These questions are formulated as point-prediction tasks, challenging the model to determine appropriate locations within free space where an object can be placed—for example, “Where on the table can I put the plate?” Spatial compatibility builds on the identified empty space to assess whether a given area can feasibly support the placement of a specific object, ensuring sufficient size and fit.
+These questions are posed in a binary format, such as “Can the chair be placed in front of the table?” Spatial configuration examines the relative spatial relationships between two objects.
+These questions use a binary format to determine whether a spatial relation holds, such as “Is the mug to the left of the laptop?”
+To enhance the model’s ability to interpret spatial instructions from different perspectives, each question-answer pair in RoboSpatial is posed from three distinct reference perspectives/frames: (a) Ego-centric from the observer’s perspective at the camera pose, (b) World-centric grounded in a global world frame, and (c) Object-centric based on a reference frame attached to the focal object.
 This multi-frame approach enables models to handle complex spatial instructions more flexibly, preparing them to better generalize to dynamic robotic contexts.
 Applying our methodology to existing indoor scene and tabletop datasets, we generate both a comprehensive training dataset and a benchmark for spatial question answering in robotics.
-RoboSpatial contains around 1M images, 5k 3D scans, and 3M annotated spatial relationships, with paired 2D egocentric images and 3D scans to make it both 2D and 3D ready, see Figure 1 for an example taken from our dataset.
-To validate the effectiveness of our dataset, we conduct comprehensive experiments with multiple state-of-the-art (SOTA) 2D and 3D VLMs.
-Results show that training VLMs on RoboSpatial significantly enhances their spatial reasoning capabilities, with models trained on our dataset outperforming existing models on our test split.
-To further demonstrate the robustness of RoboSpatial, we assess the trained VLMs on three additional downstream robotic tasks: RoboSpatial-Home, , a high-quality manually annotated indoor scene question-answering dataset, BLINK-Spatial [17], and real-world robot manipulation.
-These tasks directly evaluate the models’ ability to apply spatial understanding in practical scenarios, such as rearranging objects or responding to context-specific spatial questions within an indoor environment.
-Additionally, we test our models on an external benchmark, BLINK-Spatial [17], to measure generalization to unseen spatial reasoning tasks beyond our dataset.
-Leveraging RoboSpatial’s 3D-ready design, we also compare the performance of 2D and 3D VLMs directly.
-These results emphasize the strengths of RoboSpatial in advancing spatial understanding and reinforce the benefits of 3D-based training for enhancing the spatial reasoning capabilities of VLMs in robotic applications.
+RoboSpatial contains around 1M images, 5k 3D scans, and 3M annotated spatial relationships, with paired 2D egocentric images and 3D scans to make it both 2D- and 3D- ready.
+To validate the effectiveness of RoboSpatial, comprehensive experiments were conducted using multiple state-of-the-art (SOTA) 2D and 3D VLMs.
+Results demonstrate that models trained on RoboSpatial exhibit significantly improved spatial reasoning capabilities, consistently outperforming baseline methods on the evaluation benchmark RoboSpatial-Val, a held-out validation subset derived from the heuristically generated RoboSpatial dataset.
+To further assess the generalization and robustness of these trained VLMs, additional evaluations were conducted using three complementary benchmarks: RoboSpatial-Home, a manually collected dataset consisting of paired RGB and depth images, and two external benchmarks, BLINK-Spatial [15] and SpatialBot [2].
+These benchmarks rigorously test spatial reasoning skills in practical robotic tasks, including object rearrangement and contextual question answering in indoor environments, while also examining the models’ capacity to generalize to novel spatial reasoning scenarios beyond the original training data.
+Across all benchmarks, models trained on RoboSpatial consistently outperformed baseline methods, demonstrating the broad utility of the dataset.
+Leveraging the 3D-ready design of RoboSpatial, direct comparisons between the spatial reasoning performance of 2D and 3D VLMs were also performed.
+Although initial results indicate potential advantages for 3D models, differences in pretraining data and base LLM architectures among models render the comparison inconclusive.
+RoboSpatial is specifically designed to support both 2D and 3D research directions, enabling future studies to address these differences more conclusively.
 Our contributions are threefold:
-  • A new training dataset and benchmark, RoboSpatial and RoboSpatial-Home, comprising images and 3D scans paired with spatial questions and answers, designed with unique features including multiple reference frames, object-object spatial relationships, object-space relationships, and object compatibility.
-We plan to publicly release our data and code for generating the dataset from 3D annotated scenes.
-  • Training of VLMs on RoboSpatial, resulting in enhanced spatial reasoning capabilities that significantly outperform prior VLMs on natural language-specified robot manipulation tasks and indoor scene question answering.
+  • A new training dataset, RoboSpatial, comprising images and 3D scans paired with spatial questions and answers, accompanied by an evaluation benchmark, RoboSpatial-Val, a held-out validation set.
+Additionally, we introduce RoboSpatial-Home, a manually collected and annotated dataset designed specifically for assessing real-world spatial reasoning in indoor environments.
+These datasets uniquely incorporate multiple reference frames, object-object spatial relationships, object-space relationships, and object compatibility.
+We make the data and code for generating the dataset from 3D annotated scenes publicly available [^1] https://chanh.ee/RoboSpatial/.
+  • VLMs trained on RoboSpatial demonstrate superior spatial reasoning, outperforming SOTA baselines on language-guided robot manipulation and indoor scene question answering.
   • Comprehensive experiments assessing spatial reasoning capabilities in both 2D and 3D VLMs, comparing the difference between SOTA VLMs in real-world spatial tasks.
@@
 Recent advancements have demonstrated the potential of VLMs in various robotic applications.
-For instance, vision-and-language action models (VLAs) [27, 4, 39] enable robots to interpret and execute complex instructions and output executable robot actions.
+For instance, vision-language-action models (VLAs) [27, 65, 41] enable robots to interpret and execute complex instructions and output executable robot actions.
 Additionally, VLMs like GPT-4v [42] have been utilized for high-level task planning [55], allowing robots to generate detailed action sequences from natural language instructions.
-Furthermore, VLMs have been used for keypoint prediction [22, 16, 38], error analysis [14], grasp pose prediction [19].
+Furthermore, VLMs have been used for keypoint/mask prediction [21, 59, 40], error analysis [11, 49], grasp pose prediction [19].
 Despite these advancements, integrating VLMs [2, 6, 62] into robotic systems presents challenges.
@@
 2D image points) answers.
-From 1M images and 1.4K scans, we generate over 2M spatial question/answer pairs.
+From 1M images and 5k scans, we generate over 3M spatial question/answer pairs.
 Spatial Understanding with VLMs.
 Spatial understanding has been implicitly and explicitly part of various vision and question answering tasks [15, 1, 24, 51, 46, 28, 25, 23].
-While many benchmarks and methods have been proposed, they often come with limitations: some focus exclusively on simulations [48] or generic images [33, 41, 9, 8, 17, 26, 43], others are difficult to evaluate [48, 13], rely on complete 3D scans [58, 37, 36, 32], do not consider reference frames [58, 37, 36, 32, 8, 9, 17], or use data generated by LLMs that may lack sufficient accuracy [32].
-Furthermore, they often fail to address actionable, robotics-relevant spatial relationships such as spatial compatibility and context [13, 17, 54, 43, 26].
+While many benchmarks and methods have been proposed, they often come with limitations: some focus exclusively on simulations [53] or generic images [33, 43, 6, 5, 15, 26, 47, 44], others are difficult to evaluate due to their reliance on free-form text outputs [53, 10, 32], some rely on complete 3D scans [64, 39, 37, 32], and others do not account for reference frames [64, 39, 37, 32, 5, 6, 15, 44].
+Furthermore, many fail to address actionable, robotics-relevant spatial relationships such as spatial compatibility and context [10, 15, 58, 47, 26, 32, 44].
 Inspired by prior works on spatial reasoning [33, 26]—where the impact of reference frames and spatial configurations was explored in generic images [31, 23]—we extend spatial understanding to a robotics-specific context with actionable spatial relationships such as spatial compatibility and spatial context.
 Our aim is to enable direct application to robotic workflows, such as task planning and verification.
-To achieve this, we have developed and are planning to open-source a large-scale 2D/3D ready pretraining dataset, an automated data annotation pipeline, and trained models.
-We further show how our dataset can be used to teach spatial reasoning to a suite of vision-language models (VLMs) in in-domain and out-of-domain spatial reasoning datasets.
+To achieve this, we have developed a large-scale 2D/3D ready training dataset using our automated data generation pipeline.
+We further show how RoboSpatial can be used to teach spatial reasoning to a suite of vision-language models (VLMs) in in-domain and out-of-domain spatial reasoning datasets.
 We hope these resources lower the barrier to entry for exploring spatial understanding tailored to robotics.
 3 Approach
-In this section, we begin by explaining our selection of three spatial relationships: spatial configuration, spatial context, and spatial compatibility.
-Next, we provide an overview of the data generation pipeline used to construct RoboSpatial.
-An overview of the RoboSpatial is provided in Figure 2.
+We begin by explaining the selection of three spatial relationships: spatial context, spatial compatibility, and spatial configuration.
+Next, we describe the data generation pipeline used to construct the RoboSpatial.
+Figure 2 provides an overview of the dataset.
 3.1 Spatial Relationships
-Our dataset is organized around three core spatial relationships that we believe to address the essential aspects of spatial reasoning for robotic tasks: object configuration, object context, and object compatibility.
+The dataset is organized around three core spatial relationships that we believe address the essential aspects of spatial reasoning for robotic tasks: spatial context, spatial compatibility, and spatial configuration.
+Context allows robots to assess the relationship between objects and their surrounding space, facilitating the identification of empty or occupied areas, which is relevant for downstream applications such as path planning and obstacle avoidance.
+Compatibility focuses on whether objects can coexist or interact without conflict in a given space, which is vital for object placement, assembly, and operational safety.
 Configuration enables robots to understand and interpret the relative positioning of objects, which is crucial for directing navigation, manipulation, and interaction within complex environments.
-Context allows robots to assess the relationship between objects and their surrounding space, facilitating the identification of empty or occupied areas and aiding in efficient path planning and obstacle avoidance.
-Compatibility focuses on whether objects can coexist or interact without conflict in a given space, which is vital for object placement, assembly, and operational safety.
 Together, these spatial relationships provide a more nuanced and practical framework for robotic applications than metrics like distance—which is hard to normalize across different scales, environments, and tasks—thereby enabling robots to perform complex tasks with greater reliability.
 3.2 Dataset Generation
-The goal of the data construction pipeline is to generate a high-accuracy spatial relationship dataset with minimal human intervention by use of carefully constructed heuristics.
-Our data generation process takes as input a source scene dataset $\mathcal{D}_{s}$ that contains RGB views, camera poses, and object annotations in form of text labels and oriented 3D bounding boxes.
-The output is a dataset $\mathcal{D}$ where each datum $d_{i}=\langle I_{i},q_{i},a_{i},l_{i}\rangle$ consists of an image $I_{i}$, a question $q_{i}$, an answer $a_{i}$, and a reference frame label $l_{i}\in\{\textit{ego},\textit{world},\textit{object}\}$.
-Each question belongs to one of the three core spatial reasoning categories or a fourth category for object grounding (Table 2).
-Our data generation process consists of two stages, (1) Spatial Relation Extraction, and (2) Question-Answer Generation.
-3.2.1 Spatial Relation Extraction
-In the spatial relation extraction stage, we analyze the dataset to discover a set of spatial relations between objects or points in free space.
-Each spatial relation can be described as $\langle I_{i},a_{i},t_{i},r_{i},l_{i}\rangle$, where $I_{i}$ is an image, $a_{i}$ is an anchor object, $t_{i}$ is a target object or a target free-space point, $s_{i}$ is size measure based on an object to be manipulated, $r_{i}\in\{\textit{left},\textit{right},\textit{above},\textit{below},\textit{front},\textit{behind}\}$ is a spatial relation preposition, and $l_{i}\in\{\textit{ego},\textit{world},\textit{object}\}$ is a reference frame label.
-We use oriented 3D bounding box annotations to automatically generate spatial relationships.
-Since the 3D bounding boxes contain information about orientation (which direction the object is facing) and location in the world (including elevation), we then utilize a heuristic to automatically extract spatial relationships between objects and space.
-Spatial Configuration: Spatial configuration relationships involve determining whether one object is in a certain spatial direction relative to an anchor object.
-We first choose an anchor object and generate spatial relationships with all the objects that appear in the image.
-For each of the three reference frames, we select the direction indicated by the spatial relation.
-For example, if the spatial relation is (in front, object frame), we only select objects that are in the forward direction from the oriented bounding box.
-To avoid ambiguity, we only consider objects that appear once in the current frame and generate relationships between all object pairs in the image.
-Answers for this relationship are binary (True/False).
-Spatial Context: Spatial context relationship refers to a task where the model is required to determine points (x,y coordinates in the camera frame, also known as image space) relative to the spatial relationship with the anchor object.
-One example of such a relationship is identifying points that are in front of another object.
-Specifically, we construct a top-down map of the environment using the annotated 3D bounding boxes and randomly sample points in empty areas that are a set distance from the object.
-This distance could be based on the size of another object that is to be manipulated and placed in relation to the object.
-We apply the same reference frame logic as described for the spatial configuration heuristic.
-Answers to this relationship are numeric, i.e., a list of 2D coordinates that fit the criteria.
-Spatial Compatibility: Spatial compatibility is an extension of the spatial context relationship where the model needs to reason about whether a referenced object can fit in a region indicated relative to the anchor object.
-We use the same top-down map constructed for spatial context generation and simulate placing an object in the map using its bounding box size.
-We apply the same reference frame logic to define the area we should simulate placing an object.
-Answers for this relationship are binary (True/False).
-Question Types Category Answer type
-Spatial Relations Spatial Configuration True/False
-Spatial Context 2D points
-Spatial Compatibility True/False
-Grounding Object Referring 2D bounding box
-Table 2: Question Taxonomy.
+The goal of the data construction pipeline is to generate a large-scale, high-accuracy spatial relationship dataset with minimal human intervention, using automatic heuristics grounded in 3D geometry and 2D image views.
+The pipeline takes as input a scene dataset $\mathcal{D}_{s}$ that contains RGB images, camera poses (both extrinsic and intrinsic parameters), and oriented 3D bounding box annotations with semantic object labels.
+The output is a spatial reasoning dataset $\mathcal{D}$, where each entry $d_{i}=\langle I_{i},q_{i},a_{i},l_{i}\rangle$ consists of an image $I_{i}$, a question $q_{i}$, an answer $a_{i}$, and a reference frame label $l_{i}\in\{\textit{ego},\textit{world},\textit{object}\}$.
+Each question is derived from one of three spatial reasoning categories: spatial configuration, spatial context, or spatial compatibility.
+To support reliable object reference resolution, we also generate an auxiliary object grounding dataset that links object descriptions to 2D bounding boxes.
+To improve clarity and reproducibility, we describe the data generation pipeline in two main stages.
+We also separate the reasoning logic used for extracting 3D relationships from that used for generating 2D image-space targets.
+3.2.1 Stage 1: 3D Spatial Relation Extraction
+The first stage involves extracting spatial relationships between objects or between objects and free space, based on 3D geometry.
+Each spatial relation is defined as $s_{i}=\langle I_{i},a_{i},t_{i},r_{i},l_{i}\rangle$, where $I_{i}$ is the source image, $a_{i}$ is the anchor object, $t_{i}$ is the target object or a sampled point in free space, $r_{i}\in\{\textit{left},\textit{right},\textit{above},\textit{below},\textit{front},\textit{behind}\}$ is the relation preposition, and $l_{i}\in\{\textit{ego},\textit{world},\textit{object}\}$ denotes the reference frame.
+We use oriented 3D bounding boxes, provided by the source dataset, to compute spatial relationships.
+Each bounding box includes both the 3D location and heading of the object.
+The object’s orientation is defined by the heading vector of the bounding box, aligned with the object’s front-facing direction.
+Using this orientation, we determine the appropriate directional region (e.g., front, left) relative to the reference frame.
+For instance, a relation such as “in front of (anchor object) (object frame)” refers to the positive direction along the anchor object’s heading vector.
+These relationships are calculated independently for each of the three reference frames: the world frame is aligned with the dataset-level coordinate system; the ego frame is defined by the camera pose (i.e., camera-centered); and the object frame is defined by the local orientation of the anchor object.
+The camera extrinsics are used to transform coordinates between reference frames.
+Although the method does not require point clouds or meshes, it relies on camera intrinsics and extrinsics to project between 2D and 3D and to ensure consistent reference frame reasoning.
+For each spatial configuration task, we evaluate all visible object pairs that appear uniquely in the image, avoiding duplicate instances to minimize ambiguity.
+The resulting relationships are binary (True/False) and specify whether the spatial condition holds for the given object pair.
+3.2.2 Stage 2: 2D Spatial Point and Region Sampling
+In the second stage, we generate 2D image-space annotations for spatial context and spatial compatibility tasks.
+These rely on the 3D bounding box layout and calibrated camera parameters to map spatial relationships into image coordinates.
+For spatial context, we construct a top-down occupancy map of the scene by marking regions occupied by 3D bounding boxes.
+We then randomly sample 3D points in empty space that lie in a specified directional relation to the anchor object, following the same frame-dependent heuristics as in the configuration task.
+These points are projected into the image plane using the camera intrinsics.
+To ensure the points are valid, we filter out samples that are obstructed or occluded based on line-of-sight from the camera.
+Specifically, we perform raycasting from the camera center to each sampled 3D point, and discard points whose rays intersect any occupied bounding box volumes before reaching the target location.
+The final answer is a list of 2D $(x,y)$ image coordinates that satisfy the spatial context constraint.
+Spatial compatibility extends this idea by checking whether a target object can fit within the sampled region.
+We simulate placing a virtual bounding box, matching the size of the target object, at the candidate location on the ground plane.
+A region is considered compatible if the simulated placement does not intersect with any existing bounding boxes in the scene and provides at least a 10 cm margin along each axis.
+The simulation allows for translation and in-plane rotation of the object.
+The answer to this task is binary (True/False), indicating whether the region can accommodate the object.
 3.2.3 Question-Answer Generation
-Based on the generated relationships, we construct three types of question-answer pairs corresponding to each spatial relationship type: spatial configuration, spatial context, and spatial compatibility.
-To ensure that the model reasons with vision rather than relying on language commonsense, we structure the questions and answers to be template-based and unambiguous.
-We use a template: {object/space} {relationship} {anchor object} {reference frame}, where the spatial relationship and reference frame is defined in  Sec.
-3.2.1.
-We maintain a balanced dataset with respect to spatial relationships to avoid introducing bias into the model.
-Furthermore, we use the 2D bounding box to create an auxiliary object-referring dataset to teach models what the object is, in order to avoid errors arising from a lack of object grounding.
-Using our pipeline we generated 3M spatial relationships, which is an order of magnitude larger than other spatial reasoning datasets shown in Table 1.
-Figure 2 shows an overview of the generated question types.
+Once the spatial relations $\{s_{i}\}$ have been extracted, we generate corresponding question-answer pairs $\{d_{i}\}$ using structured templates.
+Each question follows the format: {target} {relation} {anchor} {ref.
+frame} where the relation and frame are defined in Section 3.2.1.
+To ensure that models learn from visual grounding rather than linguistic priors, we use deterministic templates that avoid ambiguity and minimize reliance on commonsense.
+Each spatial relation type—context, compatibility, and configuration—has a corresponding question format.
+Configuration and compatibility tasks result in binary (True/False) answers.
+Context questions produce a list of valid 2D coordinates in image space.
+Correctly resolving which object is being referred to in a spatial question is essential for reliable spatial understanding.
+To reduce errors arising from incorrect object identification, we additionally generate an auxiliary object grounding dataset that links object descriptions to 2D bounding boxes in the image.
+These grounding annotations are derived by projecting 3D bounding boxes into image space using camera intrinsics and extrinsics.
+This supervision helps models more accurately resolve references during spatial reasoning and is included during training.
+See Appendix B.3 for details.
+Using this pipeline, we generate around 3 million spatial relationships and their associated question-answer pairs.
+This scale is an order of magnitude larger than prior spatial reasoning datasets (see Table 1).
 4 Experiments
-4.1 Implementation
-We apply our data generation process to three scene datasets and two tabletop datasets.
-The scene datasets include ScanNet [11], Matterport3D [7], and 3RScan [52].
-The object tabletop datasets are HOPE [50] and GraspNet-1B [15].
-We retrieve the 3D bounding box annotations and embodied images from EmbodiedScan [54].
-This mix of datasets provides coverage for diverse indoor scenes featuring large objects for navigation, and smaller objects for learning priors for robot manipulation.
-In total, we generate a large scale dataset for spatial reasoning that contains around 3M spatial QA pairs with 5K 3D scans and 1M images.
-Statistics of these datasets are described in Table 3.
+4.1 Setup
+We apply the data generation pipeline to three scene datasets—ScanNet [8], Matterport3D [4], and 3RScan [56]—and two tabletop datasets—HOPE [54] and GraspNet-1B [12].
+We retrieve 3D bounding box annotations and embodied images from EmbodiedScan [58], and generate a large-scale spatial reasoning dataset covering diverse indoor environments: larger scenes for navigation and smaller object-centric setups for manipulation.
+In total, RoboSpatial includes approximately 3M spatial QA pairs across 5k 3D scans and 1M images. (Table 2 provides a breakdown; values are rounded to the nearest thousand for clarity.)
 Dataset Type Splits Images QA pairs
-Indoor Train 4914 scans 880K images 2.6M
-Validation 50 scans 1K images 3K
-Test 50 scans 1K images 3K
-Tabletop Train 150 scenes 75K images 217K
-Validation 40 scenes 3K images 3K
-Test 77 scenes 3K images 3K
+Indoor Train 4916 scans 883k images 2.7M
+Validation 40 scans 1k images 3k
+Tabletop Train 190 scenes 76k images 220k
+Validation 77 scenes 355 images 3k
 Table 2: Dataset splits for indoor and tabletop dataset. Detailed data statistics are in the Appendix.
-Figure 3: In-domain and out-of-domain results for RoboSpatial-trained models.
-Two models are shown: SL (SpaceLLaVA [8]) and RP (RoboPoint [57]), where the S- prefix and FT indicates RoboSpatial-trained models.
-For Yes/No questions, the green checkmark indicates the correct answer.
-For spatial context questions, GT indicates the correct answer.
-For out-of-domain, all images except for the top right are from RoboSpatial-Home.
 4.1.1 Trained 2D/3D VLMs
 2D VLMs.
-For our dataset, we want to compare a suite of VLMs, including ones fine-tuned for spatial understanding and ones that are not.
-We chose two base VLMs: VILA-1.5-8B [30] and LLaVA-NeXT-8B [34]; three specialized VLMs: SpaceLLaVA-13B, a community implementation of SpatialVLM [8]; RoboPoint-13B [57], a model fine-tuned on synthetic datasets to predict points in empty areas referenced by an object; and Molmo-7B [12], a new family of VLMs specializing in pointing and counting We also include a closed-source model, GPT-4o [40].
-We omit models such as SpatialRGPT [9] due to its reliance on external object mask generation as input, which omits the important aspect of object grounding.
+We evaluate several vision-language models (VLMs) using RGB-only image inputs.
+Our selected base VLMs are VILA-1.5-8B [30] and LLaVA-NeXT-8B [35].
+We also include three specialized models: SpaceLLaVA-13B (a community version of SpatialVLM [5]), RoboPoint-13B [62] (trained to predict points in empty space given an object reference), and Molmo-7B [9] (designed for pointing and counting).
+We also include GPT-4o [42] as a closed-source baseline.
+We omit models such as SpatialRGPT [6] that depend on external mask inputs, as they bypass the object grounding challenge.
 3D VLMs.
-Extending VLMs to the 3D domain introduces unique challenges due to the inherent complexity of 3D data.
-The scarcity of large-scale, high-quality 3D datasets hinders the development of end-to-end 3D VLMs.
-As such, there are few open source VLMs for 3D understanding that reason from 3D inputs.
-Among the available models, we selected: 3D-LLM [18], which lifts multi-view images into a 3D point cloud, and LEO [20], which directly takes as input 3D object point clouds by relying on a pre-processing object segmentation step.
-These models allow us to evaluate spatial understanding in 3D by processing data either through reconstructed 3D representations or directly from point clouds.
-By incorporating both approaches, we aim to assess the effectiveness of 3D VLMs in handling spatial information from different types of 3D data inputs.
-Fine-tuning. Our evaluation includes both “zero-shot” evaluation of each VLM, plus evaluation after fine-tuning on RoboSpatial for open-source VLMs. See the Appendix for further details.
-4.1.2 Spatial Understanding Evaluation
-We evaluate our method in answering 3000 questions (1000 questions per spatial relationship) from our test set, and measure how often the model produces the correct answer.
-The evaluation includes binary yes/no questions, and questions requiring numeric responses.
-For yes/no questions, we report the success rate in producing the correct response.
-For numeric questions, we report the rate at which the produced numeric answer is within the convex hull of a set of points generated by our spatial relation extraction method.
-Note that this is potentially an overly strict metric, and therefore the reported performance on numeric questions reflects a lower-bound of the true performance.
-Table 4 shows the main results.
-4.1.3 Cross-Dataset Generalization Evaluation
-Next, we evaluate the generalization capability of our method by testing it across different scene types—specifically, both indoor and tabletop scenes—to control for any bias in the annotations of the underlying datasets that make up our benchmark.
-We train on data derived from subsets of the datasets corresponding to one scene type (either indoor or tabletop) and test on held-out datasets from the other scene type, representing unseen environments.
-We expect that even when training on a subset of datasets, the performance on unseen scene types will improve if our method generalizes well.
-The results of this cross-dataset evaluation are shown in Table 5.
-4.1.4 Out-of-Domain Evaluation
-To test the out-of-domain transferability of RoboSpatial-trained models, we evaluate them on RoboSpatial-Home, our manually annotated indoor scene dataset, and the validation spatial reasoning split of the Blink [17] benchmark.
-RoboSpatial-Home is a collection of 50 crowd-sourced indoor RGBD images captured using a handheld iPhone camera equipped with a built-in depth sensor.
-Each image is manually annotated with three spatial relationships, resulting in a total of 150 questions across 50 images.
-We designed RoboSpatial-Home to test the RoboSpatial-trained model’s capability to generalize to unseen objects in real life cluttered scenes.
-Blink [17] is a visual reasoning benchmark designed to evaluate the reasoning capabilities of vision-language models (VLMs) using visual input alone.
-Blink was chosen to assess the generalization ability of RoboSpatial-trained models on unseen spatial relationships such as “next to,” “touching,” and “on top.” Unlike benchmarks such as What’s Up [26], which include spatial relationships similar to those in RoboSpatial, Blink [17] provides a more rigorous test of out-of-domain generalization.
-The validation spatial reasoning split of Blink consists of 143 binary yes/no questions about spatial configurations.
-The results of the out-of-domain generalization evaluations are presented in Table 6 and visualized in Fig. 3.
+Models operating over 3D data must handle richer, more complex spatial representations.
+We include two models that process 3D inputs: 3D-LLM [18], which reconstructs colored 3D point clouds from multi-view RGB images, and LEO [20], which operates on segmented colored point clouds of individual objects.
+These models allow us to explore spatial reasoning when models consume RGBD or point cloud representations directly.
+Fine-tuning.
+We evaluate models in both zero-shot and fine-tuned settings, using RoboSpatial to fine-tune open-source models.
+To mitigate failure cases arising from poor object grounding, we also include an auxiliary grounding dataset during training, which provides additional supervision for object reference resolution.
+This auxiliary dataset does not contribute to spatial reasoning performance.
+(See Appendix for ablation experiments.)
 Model Indoor Tabletop Average
@@
 GPT-4o [42] 63.5 25.1 59.4 62.3 27.9 66.8 49.3 52.3 50.8
-Table 4: Results of existing 2D/3D VLMs on a held-out test split of images and scans.
-All methods, for all tasks, perform better ($\uparrow$) when fine-tuned on our RoboSpatial dataset.
+Table 3: Results of existing 2D/3D VLMs on a held-out validation split (RoboSpatial-Val) of images and scans.
+All methods, for all tasks, perform better ($\uparrow$) when fine-tuned on RoboSpatial.
 The best result for each column is bolded.
-Indoor $\rightarrow$ Tabletop Tabletop $\rightarrow$ Indoor
+Model RoboSpatial-Home BLINK SpatialBench
+Configuration Context Compatibility Accuracy Accuracy
+2D VLMs
+VILA [30] 57.8 0.0 69.0 72.7 53.0
++RoboSpatial 65.9 $\uparrow$ 15.6 $\uparrow$ 78.0 $\uparrow$ 79.7 $\uparrow$ 73.6 $\uparrow$
+LLaVA-NeXT [35] 60.2 0.0 71.4 71.3 55.9
++RoboSpatial 76.4 $\uparrow$ 19.7 $\uparrow$ 80.1 $\uparrow$ 79.0 $\uparrow$ 70.6 $\uparrow$
+SpaceLLaVA [5] 61.0 0.08 34.3 76.2 47.1
++RoboSpatial 71.6 $\uparrow$ 13.1 $\uparrow$ 62.0 $\uparrow$ 81.8 $\uparrow$ 67.7 $\uparrow$
+RoboPoint [62] 59.4 2.5 80.1 63.6 44.1
++RoboSpatial 70.0 $\uparrow$ 21.3 $\uparrow$ 88.6 $\uparrow$ 70.6 $\uparrow$ 64.7 $\uparrow$
+3D VLMs
+3D-LLM [18] 39.8 0.0 35.2 N/A N/A
++RoboSpatial 55.2 $\uparrow$ 8.2 $\uparrow$ 52.3 $\uparrow$ N/A N/A
+LEO [20] 51.2 0.0 38.1 N/A N/A
++RoboSpatial 64.2 $\uparrow$ 10.0 $\uparrow$ 57.1 $\uparrow$ N/A N/A
+Not available for fine-tuning
+Molmo [9] 48.7 0.0 24.8 67.1 55.9
+GPT-4o [42] 74.0 6.6 55.2 76.2 70.6
+Table 4: Results on an out-of-domain test split comparing prior art VLMs.
+The results show improved ($\uparrow$) spatial understanding capabilities on similar domains.
+Bolded number is the best result for the column.
+4.1.2 Spatial Understanding Evaluation
+We evaluate spatial reasoning capabilities using RoboSpatial-Val, a held-out validation subset of RoboSpatial sampled from scans that are entirely unseen during training.
+This benchmark comprises 6,000 heuristically generated questions from our data generation pipeline, with 2,000 questions per spatial relation type.
+Questions fall into two categories: binary yes/no questions and coordinate prediction tasks.
+For yes/no questions, we report accuracy.
+For coordinate predictions, we evaluate whether the model’s predicted 3D location lies within the convex hull of a reference point set derived from scene geometry.
+While this convex hull criterion provides a well-defined geometric target, it is arguably overly strict—e.g., predictions near but just outside the boundary are marked incorrect.
+As a result, reported scores represent a conservative estimate of each model’s spatial understanding.
+RoboSpatial-Val serves as the primary benchmark for comparing 2D and 3D VLMs trained on RoboSpatial, enabling controlled evaluation within the same data distribution.
+Results are presented in Table 3.
+4.1.3 Cross-Dataset Generalization Evaluation
+To assess generalization across environment types, we partition the training data into indoor-scene and tabletop subsets.
+Models are trained on one type and evaluated on held-out datasets from the other.
+Despite differing object distributions and scene layouts, we observe a positive synergy between indoor and tabletop environments: training on one environment type improves spatial reasoning on the other, as shown in Table 6.
+Figure 3: In-domain (RoboSpatial-Val, top) and out-of-domain (RoboSpatial-Home, BLINK [15], middle and bottom) results for RoboSpatial-trained models.
+Two models shown: SL (SpaceLLaVA [5]) and RP (RoboPoint [62]); the -FT suffix indicates fine-tuning on RoboSpatial.
+Correct answers in green.
+All images except bottom-right in the out-of-domain rows are from RoboSpatial-Home.
+4.1.4 Out-of-Domain Evaluation
+To test the out-of-domain transferability of RoboSpatial-trained models, we evaluate on three benchmarks: RoboSpatial-Home, BLINK [15], and SpatialBench [2].
+RoboSpatial-Home contains 350 manually written spatial questions over diverse real-world RGBD scenes captured with an iPhone equipped with a depth sensor.
+We curated this benchmark to evaluate generalization to novel indoor settings with previously unseen objects.
+BLINK is a visual reasoning benchmark consisting of binary spatial questions involving relationships such as “next to,” “touching,” and “on top.” BLINK allows us to assess the ability of models to assess generalization of spatial reasoning to unseen language configurations.
+We evaluate only the spatial portion of BLINK, as that aligns with the core focus of RoboSpatial.
+SpatialBench, introduced in the SpatialBot [2] paper, uses RGB inputs and tests spatial understanding across several categories.
+We focus on the position category, which most directly aligns with our emphasis on spatial localization and placement.
+Together, these benchmarks complement RoboSpatial-Val by covering a broader range of visual and linguistic variation.
+4.2 Results
+We evaluate the effectiveness of RoboSpatial in improving spatial reasoning capabilities in VLMs across held-out and out-of-domain benchmarks.
+In this section, we focus on analyzing the model’s generalization and understanding of spatial relationships.
+We address the following questions:
+How well does RoboSpatial training generalize to unseen spatial relationships?
+Although RoboSpatial consists of template-generated QA pairs with a fixed set of spatial prepositions, we observe in Tab.
+4 that models trained on it can generalize to spatial relationships not explicitly included in the training set.
+This is particularly evident in evaluations on the BLINK dataset [15], which contains diverse prepositions such as “under,” “next to,” and “far away.” We attribute this generalization to the fact that RoboSpatial encompasses all six principal directions in 3D space (along the x, y, and z axes).
+Generalizing to new prepositions often requires mapping linguistic expressions (e.g., “on top of,” “under”) to these spatial primitives—a task at which LLMs are naturally proficient.
+For example, “on top of” often refer to “above” in a world-centric frame, while “under” maps to “below.” Moreover, prepositions such as “next to” or “beside” imply proximity between objects.
+Because RoboSpatial includes questions that require generating points near a reference object, it implicitly teaches the concept of closeness.
+This enables trained models to understand these proximity-based relationships, even if they are not explicitly represented during training.
+Do RoboSpatial-trained models understand nuanced perspectives?
+Spatial references in natural language often imply specific reference frames.
+For instance, “in front of the car” typically refers to the direction of the car’s front hood.
+In RoboSpatial-Home, we omit explicit frame specifications in the questions to evaluate whether models can align with the implicit reference frame intended by the questioner.
+We find that models trained with RoboSpatial can often infer the correct frame of reference, suggesting that they have learned to associate object geometries and orientations with spatial language.
+Figure 3 shows examples such as “Is the frame in front of the window?”, where the model accurately identifies the intended spatial relation.
+Are 3D VLMs better at learning spatial relationships than 2D VLMs?
+The findings in Tab.
+3 suggest that 3D VLMs tend to outperform 2D counterparts in spatial reasoning tasks, likely due to their ability to directly utilize depth information.
+However, this comparison is not entirely fair: models like 3D-LLM [18] and LEO [20] are pretrained on RGB-D indoor scan datasets, some of which overlap with the environments used in the source datasets (e.g., Matterport3D, ScanNet).
+This gives them prior exposure to scene geometry and object layouts, which may bias their performance.
+To support more controlled and fair comparisons in the future, we designed RoboSpatial to be compatible with both 2D and 3D modalities, allowing researchers to investigate the impact of modality, architecture, and pretraining data under unified evaluation protocols.
+Indoor Tabletop
+$\downarrow$ $\downarrow$
+Tabletop Indoor
 RoboPoint [62] 38.7 38.2
@@
 +RoboSpatial 47.2 $\uparrow$ 54.5 $\uparrow$
-[…]
-Model RoboSpatial-Home BLINK-Spatial
-Localization Affordance Compatibility Accuracy
-Open-source VLMs
-2D VLMs
-VILA [30] 53.3 12.0 52.0 72.7
-+RoboSpatial 62.0 $\uparrow$ 32.0 $\uparrow$ 58.0 $\uparrow$ 79.7 $\uparrow$
-LLaVA-Next [34] 48.0 9.3 37.3 71.3
-+RoboSpatial 58.0 $\uparrow$ 24.0 $\uparrow$ 44.0 $\uparrow$ 79.0 $\uparrow$
-SpaceLLaVA [8] 60.0 16.0 49.3 76.2
-+RoboSpatial 68.7 $\uparrow$ 38.0 $\uparrow$ 56.0 $\uparrow$ 81.8 $\uparrow$
-RoboPoint [57] 43.3 41.3 36.0 63.6
-+RoboSpatial 50.0 $\uparrow$ 54.0 $\uparrow$ 48.0 $\uparrow$ 70.6 $\uparrow$
-3D VLMs
-3D-LLM [18] 40.0 8.0 46.0 N/A
-+RoboSpatial 48.0 $\uparrow$ 36.0 $\uparrow$ 52.7 $\uparrow$ N/A
-LEO [20] 50.7 10.0 48.0 N/A
-+RoboSpatial 64.0 $\uparrow$ 40.0 $\uparrow$ 60.0 $\uparrow$ N/A
-Not available for fine-tuning
-Molmo [12] 44.7 38.0 58.0 67.1
-GPT-4o [40] 64.0 25.3 56.7 76.2
-Table 6: Results on an out-of-domain test split comparing prior art VLMs.
-The results show improved ($\uparrow$) spatial understanding capabilities on similar domains.
-Bolded number is the best result for the column.
-4.2 Results
-Our experiments on the held-out test split in Table 4, out-of-domain splits in Table 6, and real-world robot experiments in Table 7 clearly demonstrate the power of RoboSpatial in teaching spatial reasoning to VLMs where training on RoboSpatial improved spatial understanding across all spatial reasoning benchmarks.
-We further answer the following detailed questions about  RoboSpatial-trained models:
-How well does RoboSpatial training generalize to unseen spatial relationships?
-While RoboSpatial contains templated question-answer pairs and a fixed set of spatial relationship prepositions, Table 6 shows that RoboSpatial training does generalize to other spatial relationship prepositions present in Blink [17], such as “under,” “next to,” and “far away.” We highlight that RoboSpatial encompasses all six principal directions along the x, y, and z axes in 3D space, and that generalizing to novel relationships involves mapping each of these directions to one or more spatial relationship prepositions, a task at which LLMs excel.
-For example, “on top of” and “on” refer to the “above” direction in a world-centric frame, while “under” corresponds to “below.” Furthermore, we observe that prepositions such as “next to” and “beside” imply that objects need to be in close proximity to one another.
-Since RoboSpatial contains spatial context questions that teach the model to generate points close to a given object, this implicitly teaches what it means to be “close” to another object, enabling RoboSpatial-trained models to understand concepts like “next to.” However, RoboSpatial training does not seem to generalize well to questions involving human-centric perspectives, since we focus solely on objects in our dataset for robotics use cases.
-Do RoboSpatial-trained models understand nuanced perspectives?
-When humans refer to spatial relationships, we often imply a specific frame of reference.
-For example, “in front of the car” typically refers to the car’s front hood.
-Therefore, questions in RoboSpatial-Home omit explicitly specifying the reference frame to test the model’s ability to align its understanding with the questioner’s intended reference frame.
-Results in Table 6 show that models trained with RoboSpatial can automatically infer the reference frame, with Figure 3 showing examples like “Is the frame in front of the nightstand.”
-Are 3D VLMs better at learning spatial relationships than 2D VLMs?
-While our results seem to imply that 3D VLMs outperform their 2D counterparts in learning spatial relationships, we point out that 3D-LLM [18] and LEO [20] are trained with indoor scan datasets, making the comparison between 2D and 3D models not entirely fair.
-Nevertheless, we hypothesize that 3D models have an advantage in spatial understanding due to their ability to incorporate depth information.
-Therefore, we developed RoboSpatial to be both 2D and 3D ready to support further research in this area.
-Model Success Rate (%)
+Table 5: Cross-dataset generalization results between indoor and tabletop environments.
+Model Success (%)
 Open-source
@@
 GPT-4o [42] 46.9
-Table 7: Success rate comparison between robotics systems powered by zero-shot VLMs and VLMs fine-tuned on RoboSpatial.
+Table 6: Robot experiment results.
 4.3 Real Robot Experiments
@@
 The spatial- prefix indicates model trained with RoboSpatial.
-We designed a comprehensive suite of tasks that require spatial reasoning and object manipulation.
-We use a diverse set of objects, including various shapes (e.g., cubes, cylinders, boxes), colors (red, orange, blue, green), and everyday items (food items, toys) as shown in Figure 4.
-Our goal in selecting these objects is not to confound the methods with difficult-to-identify items but to make identification as simple as possible with no ambiguity.
-For each scene configuration, we asked two yes/no questions, followed by asking a pick-and-place question.
-We note that picking and placing were treated as independent tasks for evaluations.
-We conducted over 200 model queries during our experiments, where details of questions and scenes used in the experiments are in the Appendix.
-For robot experiments, we used a Kinova Jaco robot [6] paired with a ZED2 camera for RGB(D) perception and employ cuRobo [47] for collision-free motion planner.
-We evaluate the following VLMs: LLaVA-NeXT [34] and RoboPoint [57], both with and without RoboSpatial training; and Molmo [12] and GPT-4o [40] as baselines.
-Table 7 and Fig 4 presents the results of the real-world robot experiments.
-Our results show that for generic VLM such as LLaVa-Next [34], training on RoboSpatial significantly improved its spatial understanding capabilities in robot manipulation tasks.
-Furthermore, RoboSpatial-trained models were more likely to align predictions with the implied reference frame, as demonstrated in the “place in front of the pony” task, where only RoboSpatial-trained models’ predictions aligned correctly in front of the pony’s head.
-Similarly, in the “place in front of the orange juice box” task, RoboSpatial-trained models could identify an appropriate distance to place the object relative to its size.
-In contrast, for RoboPoint, the predicted point was too far to the object, making it difficult to place the object correctly.
-However, we note that GPT-4o performs comparably with RoboSpatial-trained RoboPoint.
-We attribute this to GPT-4o being able to understand much more diverse language inputs, unlike RoboSpatial, which was trained on templated language.
-Here, we show that our dataset enhances the performance of 2D VLMs in robotics tasks.
-In our robot setup, 2D VLMs project predicted points from pixel space into 3D space; however, even a 2-pixel error can translate into a displacement of 5 to 10 cm in 3D space, leading to task failure.
-Interesting future direction is exploring how perspective impacts 2D point selection in the context of VLMs.
-Another interesting future direction is to enable 3D VLMs to capably support partial 3D point cloud observations, unlike the models we currently train, which require complete colored 3D point clouds, so that they can be effectively utilized in real-world robotics scenarios.
+We design a suite of tabletop manipulation tasks requiring spatial reasoning.
+The setup includes a Kinova Jaco robot [3], paired with a ZED2 camera for RGB-D perception.
+The robot system implements actions to pick and place objects on the table using cuRobo [52] for motion planning.
+Tasks include spatial questions that require a yes/no answer, and pick-and-place instructions that require successfully controlling the robot to complete the task.
+We adopt a modular design, where the VLM is queried for spatial understanding, and the resulting predictions (e.g., target points) are passed to a separate motion planning system for execution.
+We use a range of simple, unambiguous objects—colored cubes, cylinders, food items, and toys—to ensure the challenge lies in spatial understanding rather than object recognition (Figure 4).
+In total, we conducted over 200 model queries.
+Details of the questions and scene configurations are provided in the Appendix D.5.
+We evaluate the following VLMs: LLaVA-NeXT [35] and RoboPoint [62], both with and without RoboSpatial training; and two strong baselines, Molmo [9] and GPT-4o [42].
+Table 6 and Figure 4 present the results.
+Experiments show that LLaVA-NeXT fine-tuned on RoboSpatial achieves the highest success rate across all models.
+Training with RoboSpatial enhances spatial understanding in 2D VLMs, enabling the model to correctly interpret instructions such as “place in front of the pony,” where placement is aligned with the pony’s head direction.
+It also demonstrates sensitivity to object scale, as in the task “place in front of the orange juice box,” where the model places the object at a reasonable distance.
+In contrast, baseline models such as RoboPoint frequently place objects too far from the target, likely due to limited understanding of spatial proximity.
+We also observe that spatial failures in 2D VLMs often stem from errors in projecting 2D predictions into 3D.
+Even a small 2-pixel shift in image space can translate to a 5–10 cm error in the physical world, which is significant in manipulation tasks.
+Nonetheless, models trained on RoboSpatial produce more accurate predictions, reducing these failure cases and showing the benefit of dataset-driven improvements.
+Interestingly, GPT-4o performs comparably to RoboSpatial-trained RoboPoint.
+We attribute this to GPT-4o’s broader language understanding and instruction-following ability, which partially compensates for its lack of task-specific spatial training.
+Looking forward, promising directions include investigating how viewpoint affects 2D spatial predictions, and developing 3D VLMs that can reason over partial point clouds—removing the need for complete 3D scans and making deployment in real-world systems more feasible.
 5 Conclusion
-We introduce RoboSpatial and RoboSpatial-Home, a large-scale 2D/3D spatial understanding training and evaluation dataset tailored for robotics.
+We introduce RoboSpatial, RoboSpatial-Val, and RoboSpatial-Home, a large-scale 2D/3D spatial understanding training and evaluation dataset tailored for robotics.
 Experimental results show that models trained with RoboSpatial are able to understand spatial relationships, generalize to unseen relationships, and infer nuanced reference frames, making them applicable in a wide range of tasks that require spatial understanding.
 We further demonstrate the real-world applicability of RoboSpatial with robot experiments.
-Furthermore, our automatic data generation pipeline can be used to extend the dataset whenever needed.
-We show that RoboSpatial has the potential to serve as a foundation for broader applications in robotic scenarios that require spatial understanding.
-Acknowledgement
-The authors would like to thank Youngsun Wi, Hyunho Ahn, Minjae Bae, and Hojin Yoo for providing images in the RoboSpatial-Home dataset.
-Our research is supported in part by ARL W911NF2220144 and resources from the Ohio Supercomputer Center.
-These research findings were partially derived using the Matterport dataset, the use of which is governed by its terms of service: https://kaldir.vc.in.tum.de/matterport/MP_TOS.pdf.
+In addition, our automatic data generation pipeline can be used to extend the dataset to new data sources and spatial relations.
+We show that RoboSpatial has the potential to serve as a foundation for broader applications in robotics which require spatial understanding.
+Acknowledgements
+The authors thank Youngsun Wi, Hyunho Ahn, Hojin Yoo, and Minjae Bae for providing images in the RoboSpatial-Home dataset.
+Our research was supported in part by ARL W911NF2220144 and resources from the Ohio Supercomputer Center.
+These research findings were partially derived using the Matterport dataset, the use of which is governed by: https://kaldir.vc.in.tum.de/matterport/MP_TOS.pdf.
 Appendices
 In this supplementary material, we present additional details and clarifications that are omitted in the main text due to space constraints.
-Appendix A Dataset Details
-A.1 Dataset Statistics
+Appendix A Limitations
+While RoboSpatial significantly improves spatial reasoning capabilities in VLMs, certain design choices naturally introduce trade-offs and areas for future exploration.
+First, the dataset relies on a top-down occupancy map to identify and annotate empty regions for spatial context and compatibility tasks.
+This approach simplifies reasoning about object placement on horizontal surfaces and enables efficient data generation, but it currently does not support spatial questions involving containment—such as whether an object can fit inside or under another object—which would require more detailed volumetric modeling.
+Second, although the models are deployed on a real robot using a modular approach, we do not yet explore tighter forms of integration such as training it jointly with robot trajectories [27].
+Investigating these alternatives could enhance downstream policy learning and enable more seamless end-to-end systems.
+Finally, RoboSpatial focuses on indoor and tabletop scenes containing objects commonly encountered in household environments, and does not include humans or animals.
+This reflects the nature of source datasets and our emphasis on robot object manipulation.
+While this limits coverage of social or dynamic interaction scenarios, trained models still generalizes well to out-of-distribution benchmarks like BLINK, which include humans and animals—suggesting that the learned spatial representations are broadly transferable.
+Appendix B Dataset Details
+B.1 Dataset Statistics
 We provide the full dataset statistics in Tab.
-8.
-For all training, we use only 900,000 spatial relationships, equally divided across all datasets, due to computational constraints.
+7.
+For all training, we use only 900,000 spatial relationships, sampled equally across all datasets, due to computational constraints.
 We further experiment on the effect of data scaling on Tab.
@@
 Notably, HOPE [54] and GraspNet-1B [12] contain similar tabletop images captured from different perspectives, resulting in lower dataset diversity for the tabletop environment.
-We plan to enhance the diversity of our dataset by incorporating additional tabletop datasets.
-A.2 Dataset Generation Details
-Frame Generation.
-We explain how answers are generated for each frame.
-Each question type has three answers from ego-centric, object-centric, and world-centric perspectives.
-Each frame may share the same answer since not all frames lead to unique answers.
-For our answer generation, we used the following logic:
-  • Ego-centric: The default perspective from the ego-centric view.
-  • Object-centric: Using the oriented bounding box directions, we determine the front of the object.
-With this information, we assign front, behind, left, and right.
-Above and below remain the same as in the ego-centric perspective.
-  • World-centric: Using the z-coordinate of the oriented bounding box, we modify the above and below relationships to reflect whether an object is above or below another object with respect to elevation.
-Compatibility Generation.
-For compatibility, we construct a top-down map as shown in Fig.
-5.
-Using the top-down map and the top-down 2D bounding box of the object to be placed, we determine if there exists an empty space within a threshold distance (i.e., 1% of the object’s longest width or length).
-If the object meets this condition, it is deemed compatible to be fitted.
+We plan to enhance the diversity of RoboSpatial by incorporating additional tabletop datasets.
+B.2 Choice of Spatial Relationships
+In designing the dataset, we focused on spatial relationships that directly impact robotic perception, planning, and interaction: context, compatibility, and configuration.
+These were selected to reflect the core spatial reasoning challenges that robots encounter when operating in complex, real-world environments.
+We intentionally excluded tasks such as object counting, as we consider them to fall outside the scope of spatial understanding.
+While counting is an important visual reasoning skill, it does not require reasoning about spatial relations between objects or between objects and their environment.
+For example, determining that “three cups are on the table” is a perceptual task rather than a spatial reasoning one.
+As such, counting may complement but does not substitute for the types of relational reasoning we target.
+We leave the integration of counting tasks into spatial benchmarks as future work.
+Similarly, we exclude tasks that rely solely on distance measurements.
+Although distance is a fundamental spatial quantity, it is difficult to define consistently across different environments, object scales, and robot embodiments.
+Absolute distances can vary significantly between indoor and outdoor scenes, small and large objects, or different robot perspectives, making them hard to normalize or interpret in a general way.
+Moreover, distance alone often lacks the relational semantics required for higher-level reasoning—for example, understanding that an object is behind, above, or in front of others.
+RoboSpatial instead focuses on spatial relationships that are more invariant, interpretable, and transferable across diverse robotic scenarios.
+That said, the data generation pipeline is general and could readily support auxiliary tasks involving object counting or distance estimation if desired.
+These metrics may serve as useful complements in future extensions of the benchmark or as auxiliary supervision signals in model training.
+B.3 Object Grounding Dataset
+To support accurate spatial understanding, we generate an auxiliary dataset for object grounding.
+Many spatial reasoning tasks assume that the model can correctly identify which object is being referred to in the scene.
+However, in practice, this can be a major source of error—especially in cluttered environments or when multiple instances of the similar object type are present.
+The grounding dataset provides direct supervision to help models learn to associate text descriptions with specific objects in the image.
+For each image, we include a set of object descriptions (e.g., “the keyboard” or “the chair”) paired with the corresponding 2D bounding box of the object in the image.
+These 2D boxes are projected from the annotated 3D bounding boxes using camera intrinsics and extrinsics.
+A total of 100k grounding QA pairs are generated and used during training to reduce reference ambiguity and improve object identification accuracy in spatial tasks.
+While not part of the main spatial reasoning taxonomy, grounding accuracy is a prerequisite for answering spatial questions correctly, and we find that including this data helps reduce errors caused by incorrect object identification.
+Figure 5: An example of generated top-down map of the image from 3D bounding boxes.
 Category Dataset Split Scans Images Configuration Q Context Q Compatibility Q
 Indoor Matterport3D [4] Train 1859 scans 236243 298439 298439 298439
-Validation 5 scans 100 100 100 100
-Test 5 scans 100 100 100 100
-ScanNet [11] Train 1512 scans 278402 298439 298439 298439
-Validation 12 scans 1200 500 500 500
-Test 12 scans 1200 500 500 500
-3RScan [52] Train 1543 scans 365355 298439 298439 298439
-Validation 9 scans 900 400 400 400
-Test 9 scans 900 400 400 400
-Tabletop HOPE [50] Train 50 scenes 50000 36317 36317 36317
-Validation 10 scenes 50 500 500 500
-Test 47 scenes 235 500 500 500
-GraspNet-1B [15] Train 100 scenes 25500 36317 36317 36317
+Validation 10 scans 200 200 200 200
+ScanNet [8] Train 1514 scans 280402 299039 299039 299039
+Validation 12 scans 400 400 400 400
+3RScan [56] Train 1543 scans 366755 298839 298839 298839
+Validation 18 scans 400 400 400 400
+Tabletop HOPE [54] Train 60 scenes 50050 36817 36817 36817
+Validation 47 scenes 235 500 500 500
+GraspNet-1B [12] Train 130 scenes 25620 36817 36817 36817
 Validation 30 scenes 120 500 500 500
-Test 30 scenes 120 500 500 500
 Table 7: Full dataset statistics for indoor and tabletop datasets.
-Figure 5: An example of generated top-down map of the image from 3D bounding boxes.
-Appendix B Implementation Details
-B.1 Model Training
-We further explain the training details for all 2D and 3D VLMs trained on RoboSpatial.
-For all models, we perform instruction tuning using the model weights from public repositories.
-All training is done using 8 Nvidia H100 GPUs, with the training time between 20 and 40 hours.
-VILA [30] We initialize our model from Efficient-Large-Model/Llama-3-VILA1.5-8B on Hugging Face.
-We use the fine-tuning script from the VILA GitHub repository to train our model using the default hyperparameters.
-LLaVA-NeXT [34] We initialize our model from lmms-lab/llama3-llava-next-8b on Hugging Face. We use the LLaVA-Next fine-tuning script from the LLaVA-Next repository using the default hyperparameters.
-SpaceLLaVA [8] As official code and weights for SpatialVLM [8] is not released, we use a community implementation which is endorsed by SpatialVLM [8] authors.
-We initialize our model from remyxai/SpaceLLaVA from Hugging Face.
-We use LLaVA-1.5 finetuning script from LLaVa [35] repository using the default hyperparameters.
-RoboPoint [57] We initialize our model from wentao-yuan/robopoint-v1-vicuna-v1.5-13b on Hugging Face.
-We use the fine-tuning script provided in the RoboPoint [57] GitHub repository to train our model using the default hyperparameters.
-3D-LLM [18] We initialize our model using the pretrain_blip2_sam_flant5xl_v2.pth checkpoint downloaded from the official GitHub repository.
-Since the model requires preprocessing of multiview images, we follow the author’s pipeline to process multiview images from our environments.
-Because the model does not accept image input, we append the following text in front of the question to ensure the model understands the perspective from which the question is being asked: “I am facing anchor object.” We use the default hyperparameters and train the model for 20 epochs per the author’s guidelines.
-We choose the best model based on validation accuracy.
-LEO [20] We initialize our model from the sft_noact.pth checkpoint downloaded from the official GitHub repository.
-Since LEO supports dual image and 3D point cloud input, we input both of them and modify the question as in 3D-LLM.
-We use the default hyperparameters and train the model for 10 epochs per the author’s guidelines, and choose the best model based on validation accuracy.
-We could not fine-tune Molmo [12] from allenai/Molmo-7B-D-0924 or GPT-4o [40] from the gpt-4o-2024-08-06 API due to the unavailability of the fine-tuning script at the time of this work, thus we use them as a zero-shot baselines.
-Annotation Size 100K 300K 900k (Default) 1.8M 3M (Full)
-LLaVa-Next [34] 38.1 46.7 60.5 65.8 72.4
-Table 9: Results of scaling experiment on LLaVa-Next [34] with varied spatial relationship annotations. Average accuracy on held-out test set is reported.
-B.2 Robot Setup
-For picking, we find which object the point maps to using SAM 2 [42] and execute our picking behavior on that object.
-For placing, we simply compute the 3D coordinate based on the depth value at that pixel and place the object at that coordinate.
-There were no failures due to cuRobo [47] failing.
-The experiments were purposely designed to consist of behaviors that our robot system can handle in order to avoid introducing irrelevant factors.
-The picking behavior consists of computing a top-down grasp pose and reaching it with cuRobo [47].
-To compute the grasp pose:
-  • 1.
-    We estimate the major axis of the object’s point cloud in top-down view using PCA.
-  • 2.
-    The grasp orientation is orthogonal to the major axis.
-  • 3.
-    The grasp height is based on the highest point in the object’s point cloud minus an offset of 3cm. This heuristic ensures the system can grip long objects.
-The placing behavior is the same as picking, except that an area within 5cm of the placement coordinate is used as the point cloud for estimating orientation and height, and a vertical height offset is added to account for the height at which the object was picked.
+B.4 Dataset Generation Details
+The dataset generation pipeline is detailed in the main text (subsection 3.2), which introduces a two-stage process for computing 3D spatial relationships and projecting them into 2D image space.
+Here, we expand on implementation details not covered in the main paper and provide clarification on the reasoning logic used in spatial annotation.
+Reference Frame Annotation.
+For each spatial configuration question, we label relationships from three perspectives: ego-centric (camera view), object-centric (based on object heading), and world-centric (aligned with the dataset’s global frame).
+To compute object-centric directions, we use the heading vector of each oriented 3D bounding box to define the “front” of the object.
+Left, right, behind, and front relations are then assigned accordingly.
+World-centric annotations modify vertical relationships (above/below) using global $z$-coordinates to reflect elevation.
+Surface Detection and Free Space Sampling.
+To identify support surfaces such as tables, counters, or floors, we use GPT-4o to select candidate objects that are likely to support placement.
+A top-down occupancy map is constructed from bounding boxes in the scene  Fig.
+5.
+We sample 3D points in unoccupied regions and project them into the image plane for spatial context tasks.
+Points are filtered via occlusion checks using raycasting, ensuring sampled points are visible and unobstructed.
+Compatibility Check and Object Placement.
+For spatial compatibility, we simulate placing a virtual object bounding box at candidate locations.
+The placement must fit without intersecting other objects and must allow a clearance of at least 10 cm in all axes.
+We allow in-plane rotation and translation to test flexible placement.
+This provides a binary label (True/False) indicating whether the object can be compatibly placed in the region.
+Output Format. Though RoboSpatial uses point prediction for ease of integration with robot setups, the pipeline also supports mask-based outputs and can be extended in future work.
+Appendix C Implementation Details
 Model Indoor Tabletop Average
@@
 GPT-4o [42] 52.9 38.7 56.3 62.5 30.7 63.7 49.3 52.3 50.8
-Table 10: Results of per frame accuracy of existing 2D/3D VLMs on a held-out test split of images and scans.
-All methods, for all tasks, perform better ($\uparrow$) when fine-tuned on our RoboSpatial dataset.
+Table 8: Results of per frame accuracy of existing 2D/3D VLMs on a RoboSpatial-Val.
+All methods, for all tasks, perform better ($\uparrow$) when fine-tuned on RoboSpatial.
 The best result for each column is bolded.
-Appendix C Full Results
-C.1 Data Scaling
-In Tab.
-9, we experiment with scaling the number of annotations while keeping images fixed.
-We found that even though the number of images stays consistent, increasing the number of annotations can improve performance.
-For future work, we plan to apply our data generation pipeline to a diverse set of indoor and tabletop environments to further improve the performance of our models.
-C.2 Accuracy Per Frame
+Annotation Size 100K 300K 900k (Default) 1.8M 3M (Full)
+LLaVa-Next [35] 38.1 46.7 60.5 65.8 72.4
+Table 9: Results of scaling experiment on LLaVa-Next [35] with varied number of spatial relationship annotations. Average accuracy on RoboSpatial-Val is reported.
+MMMUval MMEp MMEc MMBenchdev
+LLaVA-NeXT 39.4 1561.8 305.4 71.6
++RoboSpatial 39.8 1604.5 293.2 71.6
+Table 10: Evaluation on general-purpose multimodal benchmarks (MMMU, MME, MMBench) to assess whether training on RoboSpatial affects commonsense and factual reasoning.
+Base Auxiliary RoboSpatial Both
+LLaVA-NeXT 30.3 32.4 51.8 60.5
+Table 11: Ablation study evaluating the impact of the auxiliary grounding dataset on RoboSpatial-Val.
+C.1 Model Training
+We further explain the training details for all 2D and 3D VLMs trained on RoboSpatial.
+For all models, we perform instruction tuning using the model weights from public repositories.
+All training is done using 8 Nvidia H100 GPUs, with the training time between 20 and 40 hours.
+C.2 Model Setup
+VILA [30] We initialize the model from Efficient-Large-Model/Llama-3-VILA1.5-8B on Hugging Face.
+We use the fine-tuning script from the VILA GitHub repository to train the model using the default hyperparameters.
+LLaVA-NeXT [35] We initialize the model from lmms-lab/llama3-llava-next-8b on Hugging Face. We use the LLaVA-Next fine-tuning script from the LLaVA-Next repository using the default hyperparameters.
+SpaceLLaVA [5] As official code and weights for SpatialVLM [5] is not released, we use a community implementation which is endorsed by SpatialVLM [5] authors.
+We initialize the model from remyxai/SpaceLLaVA from Hugging Face.
+We use LLaVA-1.5 finetuning script from LLaVa [34] repository using the default hyperparameters.
+RoboPoint [62] We initialize the model from wentao-yuan/robopoint-v1-vicuna-v1.5-13b on Hugging Face.
+We use the fine-tuning script provided in the RoboPoint [62] GitHub repository to train the model using the default hyperparameters.
+3D-LLM [18] We initialize the model using the pretrain_blip2_sam_flant5xl_v2.pth checkpoint downloaded from the official GitHub repository.
+Since the model requires preprocessing of multiview images, we follow the author’s pipeline to process multiview images from the environments.
+Because the model does not accept image input, we append the following text in front of the question to ensure the model understands the perspective from which the question is being asked: “I am facing anchor object.” We use the default hyperparameters and train the model for 20 epochs per the author’s guidelines.
+We choose the best model based on validation accuracy.
+LEO [20] We initialize the model from the sft_noact.pth checkpoint downloaded from the official GitHub repository.
+Since LEO supports dual image and 3D point cloud input, we input both of them and modify the question as in 3D-LLM.
+We use the default hyperparameters and train the model for 10 epochs per the author’s guidelines, and choose the best model based on validation accuracy.
+We could not fine-tune Molmo [9] from allenai/Molmo-7B-D-0924 or GPT-4o [42] from the gpt-4o-2024-08-06 API due to the unavailability of the fine-tuning script at the time of this work, thus we use them as a zero-shot baselines.
+Appendix D More Results
+D.1 Accuracy Per Reference Frame
 We show the results per frame in Tab.
-10 for our out-of-domain test set.
+8 for RoboSpatial-Val.
 From the results, we can see a distinct difference between 2D and 3D VLMs in understanding the world-centric frame before training with RoboSpatial.
@@
 However, we hypothesize that, due to their design—specifically, the lack of a means to visually inject perspective information since they require complete 3D point clouds—3D VLMs still lag behind 2D VLMs on ego-centric and object-centric frames.
-C.3 Robot Experiments
-We present additional results from our robot experiments in Fig.
+D.2 Data Scaling
+In Tab.
+9, we experiment with scaling the number of annotations while keeping images fixed.
+We found that even though the number of images stays consistent, increasing the number of annotations can improve performance.
+For future work, we plan to apply the data generation pipeline to a diverse set of indoor and tabletop environments to further improve the performance of the models.
+D.3 Commonsense Knowledge Retention
+To ensure that training on RoboSpatial does not degrade a model’s general reasoning or commonsense capabilities, we evaluate the RoboSpatial-trained model on a suite of standard multimodal benchmarks: MMMU [63], MME [14], and MMBench [36].
+As shown in Table 10, the RoboSpatial-trained model maintains or slightly improves performance across all benchmarks, suggesting that spatial fine-tuning preserves broader knowledge capabilities.
+D.4 Ablation of the Auxiliary Grounding Dataset
+As shown in Table 11, training on the auxiliary dataset alone yields a small improvement over the base model (+2.1), but it falls far short of the gains achieved with RoboSpatial, which is explicitly designed to teach spatial reasoning.
+This confirms that grounding supervision alone is insufficient for spatial understanding.
+However, combining both datasets leads to the best performance, suggesting that improving object localization can complement spatial supervision when jointly trained.
+D.5 Robot Experiments Details
+D.5.1 Robot Setup
+For picking, we find which object the point maps to using SAM 2 [45] and execute the picking behavior on that object.
+For placing, we simply compute the 3D coordinate based on the depth value at that pixel and place the object at that coordinate.
+There were no failures due to cuRobo [52] failing.
+The experiments were purposely designed to consist of behaviors that our robot system can handle in order to avoid introducing irrelevant factors.
+The picking behavior consists of computing a top-down grasp pose and reaching it with cuRobo [52].
+To compute the grasp pose:
+  1. 1.
+    We estimate the major axis of the object’s point cloud in top-down view using PCA.
+  2. 2.
+    The grasp orientation is orthogonal to the major axis.
+  3. 3.
+    The grasp height is based on the highest point in the object’s point cloud minus an offset of 3cm. This heuristic ensures the system can grip long objects.
+The placing behavior is the same as picking, except that an area within 5cm of the placement coordinate is used as the point cloud for estimating orientation and height, and a vertical height offset is added to account for the height at which the object was picked.
+D.5.2 Additional Results
+We present additional results from the robot experiments in Fig.
 6.
@@
 Question: pick lone object
+LLaVa-Next [35] ×
+LLaVa-Next-FT [35]
+RoboPoint [62] ×
+RoboPoint-FT [62]
+Molmo [9]
 GPT-4o [42] ×
+Question: Is there room to slot the pancake mix in the middle of the row of boxes
+LLaVa-Next [35]
+LLaVa-Next-FT [35]
+RoboPoint [62] ×
+RoboPoint-FT [62]
 Molmo [9]
-LLaVa-Next [34] ×
-S-LLaVa-Next [34]
-RoboPoint [57] ×
-S-RoboPoint [57]
-Question: Is there room to slot the pancake mix in the middle of the row of boxes
 GPT-4o [42]
-Molmo [12]
-LLaVa-Next [34]
-S-LLaVa-Next [34]
-RoboPoint [57] ×
-S-RoboPoint [57]
 Question: Is there space in the white container for the orange juice box
 LLaVa-Next [35] ×
-S-LLaVa-Next [34]
+LLaVa-Next-FT [35]
 RoboPoint [62] ×
-S-RoboPoint [57] ×
+RoboPoint-FT [62] ×
 Molmo [9] ×
@@
 LLaVa-Next [35] ×
-S-LLaVa-Next [34]
+LLaVa-Next-FT [35]
 RoboPoint [62]
-S-RoboPoint [57] ×
+RoboPoint-FT [62] ×
 Molmo [9] ×
@@
 LLaVa-Next [35] ×
-S-LLaVa-Next [34]
+LLaVa-Next-FT [35]
 RoboPoint [62]
-S-RoboPoint [57]
+RoboPoint-FT [62]
 Molmo [9] ×
@@
 LLaVa-Next [35]
-S-LLaVa-Next [34] ×
+LLaVa-Next-FT [35] ×
 RoboPoint [62]
-S-RoboPoint [57]
+RoboPoint-FT [62]
 Molmo [9] ×
@@
 LLaVa-Next [35] ×
-S-LLaVa-Next [34]
+LLaVa-Next-FT [35]
 RoboPoint [62]
-S-RoboPoint [57]
+RoboPoint-FT [62]
 Molmo [9]
@@
 LLaVa-Next [35] ×
-S-LLaVa-Next [34]
+LLaVa-Next-FT [35]
 RoboPoint [62]
-S-RoboPoint [57]
+RoboPoint-FT [62]
 Molmo [9]
@@
 LLaVa-Next [35]
-S-LLaVa-Next [34]
+LLaVa-Next-FT [35]
 RoboPoint [62] ×
-S-RoboPoint [57] ×
+RoboPoint-FT [62] ×
 Molmo [9]
@@
 LLaVa-Next [35] ×
-S-LLaVa-Next [34] ×
+LLaVa-Next-FT [35] ×
 RoboPoint [62]
-S-RoboPoint [57]
+RoboPoint-FT [62]
 Molmo [9] ×
@@
 Question: place on the object to the left of macaroni and cheese
+LLaVa-Next [35] ×
+LLaVa-Next-FT [35]
+RoboPoint [62]
+RoboPoint-FT [62]
+Molmo [9]
 GPT-4o [42] ×
+Question: is there an object that is not in a stack?
+LLaVa-Next [35]
+LLaVa-Next-FT [35]
+RoboPoint [62]
+RoboPoint-FT [62]
 Molmo [9]
+GPT-4o [42]
+Question: is there space to place one of the cans on the cheez-it box?
 LLaVa-Next [35] ×
-S-LLaVa-Next [34]
-RoboPoint [57]
-S-RoboPoint [57]
-Question: is there an object that is not in a stack?
-GPT-4o [40]
-Molmo [12]
-LLaVa-Next [34]
-S-LLaVa-Next [34]
-RoboPoint [57]
-S-RoboPoint [57]
-Question: is there space to place one of the cans on the cheez-it box?
+LLaVa-Next-FT [35] ×
+RoboPoint [62] ×
+RoboPoint-FT [62] ×
+Molmo [9] ×
 GPT-4o [42] ×
+Question: pick the highest object on the stack of two objects
+LLaVa-Next [35] ×
+LLaVa-Next-FT [35] ×
+RoboPoint [62] ×
+RoboPoint-FT [62] ×
 Molmo [9] ×
-LLaVa-Next [34] ×
-S-LLaVa-Next [34] ×
-RoboPoint [57] ×
-S-RoboPoint [57] ×
-Question: pick the highest object on the stack of two objects
 GPT-4o [42] ×
-Molmo [12] ×
-LLaVa-Next [34] ×
-S-LLaVa-Next [34] ×
-RoboPoint [57] ×
-S-RoboPoint [57] ×
 Figure 6: Additional robot experiments.
 A green check mark indicates that the model answered correctly.
-The S- prefix denotes a model trained with RoboSpatial.
+The -FT suffix denotes a model trained with RoboSpatial.
 The questions are purposely not cleaned to reflect realistic language inputs.
-Appendix D More Qualitative Examples
+D.6 More Qualitative Examples
 Fig.
-7 and Fig.
 7 present additional qualitative comparisons between models trained on RoboSpatial.
-Our findings demonstrate that models trained on RoboSpatial consistently exhibit spatial understanding in the challenging RoboSpatial-Home dataset, even outperforming closed models like GPT-4o [40].
+The findings demonstrate that models trained on RoboSpatial consistently exhibit spatial understanding in the challenging RoboSpatial-Home dataset, even outperforming closed models like GPT-4o [42].
 However, we observed that object grounding is a crucial prerequisite for spatial understanding; the improvement is often hindered by the model’s inability to ground objects in cluttered scenes, where GPT-4o performs more effectively.
-Additionally, in Fig.
-8, we show that the RoboSpatial-trained model successfully generalizes to unseen spatial relationships in Blink-Spatial [17], including those involving distance, such as ”touching.”
-Figure 7: Qualitative results on RoboSpatial-Home.
-For spatial context questions, we omit the baseline models which fail to generate a point inside the image.
-Furthermore, the first predict point for each model is shown in the image.
-Figure 8: Qualitative results on Blink-Spatial [17]. RoboSpatial-trained model can generalize to unseen spatial relationships.
+Additionally, we show that the RoboSpatial-trained model successfully generalizes to unseen spatial relationships in BLINK-Spatial [15], including those involving distance, such as ”touching.”
+Question: Pinpoint several points within the vacant space situated to the left of the pot.
+Answer
+LLaVa-Next [35]
+LLaVa-Next-FT [35]
+RoboPoint [62]
+RoboPoint-FT [62]
+Molmo [9]
+GPT-4o [42]
+Question: Pinpoint several points within the vacant space situated behind the trash bin.
+Answer
+LLaVa-Next [35]
+LLaVa-Next-FT [35]
+RoboPoint [62]
+RoboPoint-FT [62]
+Molmo [9]
+GPT-4o [42]
+Question: Can the lamp fit in front of the shelf?
+Answer Yes
+LLaVa-Next [35] ×
+LLaVa-Next-FT [35]
+RoboPoint [62] ×
+RoboPoint-FT [62]
+Molmo [9] ×
+GPT-4o [42] ×
+Question: Is the lamp above the shelf?
+Answer Yes
+LLaVa-Next [35] ×
+LLaVa-Next-FT [35]
+RoboPoint [62] ×
+RoboPoint-FT [62]
+Molmo [9] ×
+GPT-4o [42]
+Question: Is the dining table touching the donut?
+Answer Yes
+LLaVa-Next [35] ×
+LLaVa-Next-FT [35]
+RoboPoint [62] ×
+RoboPoint-FT [62]
+Molmo [9] ×
+GPT-4o [42] ×
+Question: Can the pot fit above the fridge?
+Answer Yes
+LLaVa-Next [35] ×
+LLaVa-Next-FT [35]
+RoboPoint [62] ×
+RoboPoint-FT [62]
+Molmo [9] ×
+GPT-4o [42] ×
+Question: Is the chair behind the bed?
+Answer Yes
+LLaVa-Next [35] ×
+LLaVa-Next-FT [35]
+RoboPoint [62] ×
+RoboPoint-FT [62]
+Molmo [9] ×
+GPT-4o [42] ×
+Question: Is the couch under the suitcase?
+Answer Yes
+LLaVa-Next [35] ×
+LLaVa-Next-FT [35]
+RoboPoint [62] ×
+RoboPoint-FT [62]
+Molmo [9] ×
+GPT-4o [42] ×
+Figure 7: Qualitative results on spatial reasoning benchmarks.
+The -FT suffix denotes a model trained with RoboSpatial.
+The first three rows show examples from RoboSpatial-Home, covering spatial context, spatial compatibility, and spatial configuration.
+For spatial context questions, only the first predicted point from each model is shown.
+The fourth row shows generalization to unseen spatial relationships on the Blink-Spatial [15] dataset, demonstrating that the RoboSpatial-trained model can transfer to unseen relationships.